A Complete Bibliography of Publications in Information Processing Letters: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 November 2024 Version 1.13

Title word cross-reference

(Δ + 1) [FPP23]. (l, r) [JJD22]. (s, d) [GPWM24]. 2 [BN22, BS21, Dra20, DNS20, JWZ24, JKL21, MM20, OFA21, PW21, Ray24, Sok20, Zam22]. 22k [Nut25]. 3 [PCC20, WL22, ZC23]. 3Δ − 1 [LLLW23]. 4/3 [GL25]. A^{⋆} [DBRB21]. D [LMT25, BKK23, JJD22, vBS20]. ϵ [AY21]. F [GPWM24, LS23]. γ [BDH21, CX21]. k [BFJ22, BHMP22, CFHH21, DLN^{+}23, DH23, GOR^{+}22, HK20, LT25, Nut25, Sok20, UW21, Yan23]. K_{t} [Tsu21c]. l [HJHZ22]. m [JJD22, ZZLC22]. n [Yan23]. P_{3} [CLWX25]. P_{≥3} [GW21]. P_{k} [CLWX25]. P_{t} [BHMP22, Feg23]. Π_{2} [Man21]. pq [YKC24]. R [CWW20]. R_{g} [YQL22]. st [Nut25]. y [Mas21].

-approximating [Dra20]. -approximation [Nut25]. -ary [Yan23]. -center [HK20]. -CIST [PCC20]. -coloring [FPP23]. -colouring [BHMP22]. -Complete [BDH21]. -conditional [YQL22]. -connected [JKL21]. -connectivity [ZC23]. -copies [AY21]. -critical [CX21]. -cubes [Yan23]. -cuts [Zam22]. -Dimensional [Sok20, BKK23, DNS20, Ray24, WL22]. -Disjoint [JWZ24]. -distance [JJD22]. -domination [JJD22]. -edge-connected [BN22]. -factor [GW21]. -free [BHMP22, Feg23, Tsu21c]. -hard [Man21]. -Hitting [vBS20]. -hypergraphs [GPWM24]. -hypernetworks [GPWM24]. -isogenous [HJHZ22]. -isolation [CLWX25]. -labeling [DLN^{+}23]. -length [LMT25]. -matchings [PW21]. -means [BFJ22, GOR^{+}22]. -means/median [BFJ22]. -monotone [Mas21]. -paths [Nut25]. -planarity [UW21]. -Poset [OFA21]. -restriction [LT25]. -Steiner [ZZLC22]. -systems [LS23]. -Task [MM20]. -trees [BS21]. -tuple [JJD22]. -uniform [CFHH21].

1.5D [KS23]. 158 [RT23]. 174 [BH22a].

2020 [Ano20a, Ano20b, Ano20n, Ano20o, Ano20q, Ano20p, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v]. 2021 [Ano21a, Ano21b, Ano21c, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p]. 2022 [Ano22a, Ano22b, Ano22i, Ano22j, Ano22k, Ano22l]. 2023 [Ano23a, Ano23f, Ano23g, Ano23h]. 2024 [Ano24a, Ano24f, Ano24g, Ano24h]. 2025 [Ano25c, Ano25d].

3- [CX21].

ability [APZT22]. abstract [TZZ25, WL21]. abstraction [TGS25]. accepted [PS20]. access [EK20]. accuracy [AM20]. actively [KP21]. acyclic [EHL^{+}21, Kam23]. adaptive [BEL20, PH23]. additive [Sev20]. adversarial [PR24]. adversaries [BEL20]. against [BEL20, ID23, MSYY24]. agent [AF20]. agents [LZG22]. Aho [LP22]. Airports [JN25]. algebraic [Bra22]. Algorithm [HPP20, SPG22, AK22a, ABM20, BMWW22, BN22, DGI21, DGM25, DBRB21, DFW22, EA24, FS20, FPP23, FKMS20, KK21, Lou20, LKC22, Nut25, Ohs21, OFA21, PW21, RT20a, Tsu20b, Tsu21b, Tsu23a, Tsu23b, XK22, ZZLC22, Zsc22, vIKMN22]. Algorithmic [CLWX25, PF23]. Algorithms [APS25, CCJS22, RRS25, Tsu21a, WXC24, ABT21, BEL20, DFL^{+}20, Doe21, FGIK24, FHL21, GTZ25, GLW23, HR20, Ito25, Lev22, RV23, Tsu20a, YL22]. Alignment [SPG22]. All-pairs [LP22]. alliances [GM22]. allocation [GLW23, HW21]. allowing [LL24b]. almost [AY21, ZKP^{+}24]. alphabets [BC21, LdOOW24]. alternating [PCC20]. amid [AS21]. analyses [DGM25]. analysis [DGI21, Doe21, JA20, MSS24, WZDZ22]. Analyzing [BCKP23]. annihilation [RR23]. Annotated [Sak20]. anonymity [PCO20]. anonymous [RT20b, RT23]. answer [LPT20]. Anti [BCEM24]. antipowers [FRS20]. application [MSYY24]. applications [Dür23, MWN^{+}22, WZDZ22]. approach [KL20]. approximate [AM20, DFLS23, GKNS23, Man21, Mir24]. Approximating [Zim22, Dra20, PH23]. Approximation [GJ23, GLW23, HR20, Ism24, RT20a, YL22, BN22, FS20, FGIK24, GOR^{+}22, LW23, MS20a, Nut25]. approximations [CM22, Fuj23]. April [Ano20a, Ano21a, Ano22a]. APTAS [Ray24]. arbitrary [PP24]. arc [HS24a]. arc-connectivity [HS24a]. arcs [HS24a]. Ardila [Sha21]. area [RE20, vdHKL^{+}20]. argumentation [ENRV23, TZZ25]. arguments [Doe21]. arithmetic [Mac24]. arrangement [APEiC24]. arrangements [Sax21]. array [BIM21, Lou20]. ary [Yan23]. aspects [CLWX25]. assignment [MSYY24]. asymptotically [ZZLC22]. asynchronous [Bha22]. attack [TP24, ZCWW21]. attacks [ID23, SI22b, ZY23]. augmented [PZ24]. August [Ano21b, Ano22b, Ano23a, Ano24a]. authenticity [LL24b]. autocorrelation [YKC24]. automata [BH22a, DSTZ24, Doy25, FQSW20, IK22, KP21, PS20, BH22b]. automatic [CIM20, NS24]. average [Sup22, TF23]. avoidance [CSS23]. axiom [Che23]. axiomatization [WL23]. axis [AS21, AA24]. axis-parallel [AS21, AA24]. bad [HK20]. balanced [Ito25, LL25]. balancing [HW22]. balls [Aba21]. Banerjee [KK21]. base [Sup22]. based [GB21, GKP22, HJHZ22, HvIMS25, LTT23, PK24, SP20, TP24, TZZ25, ZXY^{+}22]. bases [WL21]. batch [LG23, WTP24]. batches [JZ22, JZ23]. BC [Pou22]. BC- [Pou22]. be [KN20]. become [MPS22]. bends [LMO^{+}22]. bent [LPT20]. Berstel [Sha21]. between [AM20, Jai20, MG20]. beyond [Cha25]. bicluster [XK22, Tsu21b]. bicolored [AABC20]. Bicriteria [LG23]. bicritical [CX21]. Bijections [ENRV23]. bin [JZ22, JZ23, Ray24]. binary [DFW22, GHKY20, LdOOW24]. binding [GW21]. bipartite [HW22, MSHS23, Sch24, TV23, VP20, ZWWC22]. birthday [WZDZ22]. bisimilarity [CT20]. bisimulation [WL23, WYL24]. bit [ID23]. bit-parallel [ID23]. blind [SP20, Rab22]. block [UNSI24]. blockchain [LLP20]. blocks [BKS23]. Blue [AA24]. Board [Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano23b, Ano23c, Ano23d, Ano23e, Ano24b, Ano24c, Ano24d, Ano24e, Ano25a, Ano25b]. Bondy [AK22b]. Boolean [GRZ24]. border [IS22]. bordered [BIK23, Gab24]. Borders [GS21]. Bottleneck [BMS20]. bound [BKK23, CLWX25, CEP25, EHL^{+}21, GW21, GHN^{+}25, GL25, LXZW23, Man25, MS20b, Ruk20, Sah22a, Sup22, WZ25]. bounded [Chi20, HHT23, Kno21, SS22]. Bounds [RV23, CHTW21, Chi20, DG23, DGM25, Dür23, FRS20, GH25, GKP22, HKP21, YQL22]. boxes [AS21]. Braess [DFLS23]. Branching [WYL24, CT20, Ito25]. break [LL24b]. Bruijn [BC21]. Bubble [KM21]. Bubble-sort [KM21]. Büchi [Goe20]. buttons [Tsu20b]. byte [UNSI24]. byte-wise [UNSI24].

C [BRS21]. C-semirings [BRS21]. cactus [Fri21, HHMM20, Tsu23a]. cactuses [vIMM23]. cakes [TF23]. calculus [Fio22, Mac24]. can [JS20]. cancer [HR20]. capacitated [FS20]. Caratheodory [DFLS23]. cardinality [WXC24]. Caristi [Ish21]. case [MSS24]. Catalan [BCKV21]. CCA [HYZ^{+}20]. CCA-secure [HYZ^{+}20]. cellular [WYZ^{+}24]. center [EA24, HK20, MMCH20, MS24a]. certificate [PK24, ZXY^{+}22]. certificate-based [PK24, ZXY^{+}22]. chains [Vah24]. changeable [MMHX20]. channel [ID23]. characterization [CX21, RRS20]. Checking [MS24b, Moo22]. Chernoff [DG23, DGM25]. Chinese [LKC22]. chordal [AK22a, Dra20]. chromatic [Jac21, Sev20, TY23]. cipher [UNSI24]. circle [GKL^{+}23]. circuits [AT24, DKMS24]. CIST [PCC20]. class [EAE21, LS23, MY18, NS24, YKC24, ZKP^{+}24]. classes [APS25, CCJS22, MG20]. Claus [JR20]. clausal [Fio22]. clique [Liu23]. cliques [IK22]. closed [Ruk20]. Cluster [Tsu22]. clustering [Den22, Lab24, GJ23]. co [PRM24, Tsu23b]. Co-Path [Tsu23b]. co-secure [PRM24]. cocliques [IK22]. code [TP24]. code-based [TP24]. codes [BKS23, GHKY20, GK23, HS20, Pud22, RV23, WL22]. coffin [Vol23]. Coffman [RT20a]. cograph [Tsu20a]. cographs [KK21]. coin [DHP^{+}22]. collector [Sch21]. collision [Aba21]. colored [JR23]. coloring [DNS20, FHL21, FPP23, Sah22a, VP20]. colorings [LLLW23]. colors [LLLW23]. colouring [BHMP22]. combination [LZG22]. combinatorial [MT21, PW21]. Combining [CIM20]. Comment [LKC22]. commodities [FGS23]. common [Bli20, DBRB21, KHO21]. communicabilities [GH25]. communication [DHP^{+}22]. compact [KP21]. Comparing [GKNS23, CST22]. compatibility [HPR22, XN20]. compatible [OT21]. competitive [DGI21]. complementary [HS20]. complementation [APS25, Doy25]. complementing [IK22]. complements [VP20]. Complete [BDH21, AFK^{+}24, Bed21, HW22, MT20, UW21, ZQKL24]. completeness [GM24, IS22, Vah24]. completion [FKMS20]. complexities [PH23]. Complexity [CST23, FHL21, AY21, Ami21, Cha25, CST22, DHW22, DHP^{+}22, IS22, Ish21, KS20, Lab24, LF20, Mol22, MK20, PRM24, RSRM23, RRS25, Sel24, YKC24]. compressed [VDPT25]. compression [CWW20]. Computability [Eng21]. computable [Dra20]. computational [Lab24, LF20]. Computing [DH23, HT21, Klu24, AM20, AY21, BCV21, Lou20, Man24, Pou22]. concentration [MS20b]. concerning [Kos23]. concrete [AC21]. condition [JKL21, WQ20]. conditional [YQL22]. confidentiality [LL24b]. configuration [JR20]. conflict [RRS25]. conflict-free [RRS25]. conjecture [BS23, RT20a]. conjectures [Sch21]. conjunctive [tCFJL24]. connected [BN22, HKR21, HR20, JKL21, PP24, Ste20, WXC24]. connectedness [An22]. Connectivity [ZC23, GRZ24, HS24a, HS24b, LLC21, Yan23]. coNP [AFK^{+}24]. coNP-complete [AFK^{+}24]. Conquer [SPG22]. consensus [HHT22]. considering [ZXH20]. Constant [CDP23, DFW22, FS20, LXZW23]. constant-time [DFW22]. Constrained [Goe20, AR22, DBRB21, LF20]. constraint [ABM20, LL24a]. constraints [BFM23, CST23, HW21, Mol22, YL22]. construction [HYZ^{+}20]. constructions [GHKY20]. constructive [MS20b]. consumption [FHL^{+}23, FKMS20]. context [Kos25]. context-free [Kos25]. continued [PB23]. continuous [Mir24, WL21, ZXZ^{+}23]. control [VDPT25]. controller [FHL^{+}23]. convergence [CEP25]. convex [Bae22, BCK23b, GGSdS20, MRD25, vdHKL^{+}20]. convexity [ACG^{+}24]. copies [AY21]. Corasick [LP22]. Correct [EAE21]. Correcting [KK21, WW24, APZT22]. correctness [Sut20]. Corrigendum [BH22a, RT23]. corruption [Alw20]. cost [BFJ22, HW21, LG23, ZXH20]. costs [BRS21]. Count [VPD24, VPT24]. Counter [TGS25, HKP21]. Counting [DFMHVHT21, Bae22, Ohs21, QW24]. coupling [JPV22]. coupon [Sch21]. Cover [OFA21, Kno21, PH23, AA24]. coverability [EHL^{+}21]. coverage [HR20]. covered [Tan22]. Covering [BDBK^{+}25, GGSdS20, FGIK24]. covers [RRS20]. covet [TF23]. cow [BKK23]. cow-path [BKK23]. critical [CX21]. crossing [OT21]. crossings [DFMHVHT21]. Cryptanalysis [LTT23, OPD23, ZXY^{+}22, ZCWW21]. cryptographic [PCO20]. cryptography [HJHZ22]. cryptosystem [LKC22]. CSPs [Sta22]. cube [Ste20, ZCWW21]. cube-attack-like [ZCWW21]. cube-connected [Ste20]. cubes [JZ23, Ste20, Yan23, ZC23]. cuckoo [MP23]. curve [HJHZ22]. cut [Feg23, JKL21]. cuts [RRS25, Zam22]. cycle [ACG^{+}24, EA24]. cycles [BMS20, DE23, DS21, KM21, LL25, Ste20, Tan22, WQ20]. Cyclic [KLM23].

dark [ACG23]. data

[Gia21, MMCH20, WHN25, YL22]. deadlines [Sin23]. December [Ano20b, Ano21c]. decidability [Kos23]. Deciding [MS23, BMWW22]. decision [DKMS24, FHL^{+}23]. decomposition [Hua23]. defensive [GM22]. definability [FQSW20]. definitions [SC22]. Defying [Cha25]. degree [An22, Chi20, Sha23]. degrees [Zam22]. delay [WXC24]. deleted [Zam22]. Deletion [Tsu21c, Tsu21a, Tsu22, Tsu24, Tsu23a]. demand [JN25]. demands [FS20]. dense [BCD20]. densest [DH23]. Density [PK23]. dependent [BCV21, YHK24, Zei23]. depreciable [ZXH20]. depth [Chi20, DKMS24, LM22]. depth-2 [DKMS24]. derangements [MT23]. Design [MS20a]. despite [GKL^{+}23]. detecting [KS20]. detection [Aba21, Alw20, Bra22]. detectors [Mil21]. Determining [Bha22, PB23]. Deterministic [PP24, ABM20, DSTZ24, HYZ^{+}20, Kos25, MS24b, Tsu23a, Tsu23b]. detours [JWZ24]. diagnosability [YQL22]. diagonal [DK21]. diameter [AK22b]. diameter-revealing [AK22b]. difference [CST23]. differential [GKNS23]. differentially [Man25]. Differentiators [Mil21]. diffusion [SI22a]. digraph [Ohs21]. digraphs [Xia20]. dimension [CEP25, Man24, Sax21, ZY23]. Dimensional [Sok20, AHKBS22, BKK23, DNS20, Gia21, Ray24, WL22]. dimensions [Mir24]. directed [BS21, Fuj23, GB21, JWZ24]. discounts [Den22]. discovery [HR20]. discrepancy [Man21]. Discrete [HW22, MS24a, GM24, WZDZ22, WZ25]. Disjoint [JWZ24, DS21, LL25, Nut25, WQ20]. Dispersion [MRD25, Sha20]. distance [An22, Bod22, BCK23b, Gia21, JJD22, KLM23, LL24a, Mir24, Sha20, WY20]. distance-preserving [Bod22]. distances [FK25]. distinct [PK23]. Distinguisher [CWW20]. distributed [BEL20, DFL^{+}20, FPP23, GB21, Liu23]. distribution [BMWW22, DFW22, Sch21]. Distributivity [GS22]. Divide [SPG22, ZC23]. Divide-and-Conquer [SPG22]. divide-and-swap [ZC23]. division [Suk25]. document [Lou20]. Domain [ID23]. Domain-oriented [ID23]. dominated [AFK^{+}24]. dominating [AK22a, BN22, Fuj23, PRM24, PF23]. Domination [MP20, HPP20, JJD22, KK21, LMMZ20, Sah22b]. double [Sup22]. double-base [Sup22]. down [Doy25, MS24b]. downcast [Moo22]. Drawing [Mas21]. drawings [Bie22, HMR24, RE20, Sch21]. driven [Sak20]. driver [HR20]. drones [BDBK^{+}25]. dual [GHKY20, HS20, MSYY24]. Dynamic [DKP^{+}20, JA20, JS20, KN20]. dynamics [JPV22].

eager [KN20]. easy [CM22, MPS22]. eccentric [Pou22]. eccentricity [Dra20]. Edge [Tsu21c, BN22, DFL^{+}20, FHL21, Fuj23, LLLW23, Nut25, Tsu20a, Yan23]. edge-coloring [FHL21]. edge-colorings [LLLW23]. edges [DS21, FGIK24, LL25, PW21]. editing [XK22, Tsu21b]. Editorial [Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano23b, Ano23c, Ano23d, Ano23e, Ano24b, Ano24c, Ano24d, Ano24e, Ano25a, Ano25b]. effect [PR24]. Effective [FQSW20, WHN25]. effectively [BY24]. effectiveness [CDDN21]. effects [YHK24]. efficiency [CHTW21]. Efficient [MT23, GK23, ZXY^{+}22, tCFJL24]. Efficiently [RR23]. Egalitarian [Sut20]. Egerváry [RRS20]. election [SC22]. electrical [BMWW22]. elements [MW23]. Embedded [Yan23, BRS21]. embeddings [LMO^{+}22]. empty [Bae22]. encryption [HYZ^{+}20, ZXZ^{+}23]. End [ZWWC22]. energy [FKMS20]. enhanced [LKC22]. Entailment [EIP22]. entropy [GKP22, Sah22b]. enumerating [WXC24]. enumeration [BDK^{+}24]. environments [PK24, WYZ^{+}24, ZXY^{+}22]. envy [Kam21]. envy-free [Kam21]. equal [RR23]. equalized [PCC20]. equilibria [DH24, Goe20]. equipment [ZXH20]. equivalence [Bha22, CT20, MS23]. equivalent [CGG^{+}23b]. Erasure [APZT22]. Eratosthenes [HS25]. ergodic [Vah24]. error [APZT22, DHP^{+}22, Ism24]. Escaping [DFLS23]. ESRPKC [LKC22]. Established [EIP22]. estimation [VPT24]. Euclidean [AHKBS22, Gia21, GOR^{+}22]. Eulerian [AT24]. evaluation [HJHZ22]. Even [SI22b, JKL21]. Even-Mansour [SI22b]. evolutionary [Doe21]. Exact [HHT23, AK22a]. examples [LdOOW24]. exchange [OT21]. exclusion [RT20b, RT23]. existence [Goe20, WQ20]. expanding [GM24]. expected [DKP^{+}20]. explainable [Lab24]. exploration [Fri21, WYZ^{+}24]. Exploring [UNSI24]. exponentially [JS20]. Exposure [RSRM23]. expressions [Sak20]. Expressive [GM24]. EXPTIME [MT20, Bed21]. EXPTIME-complete [MT20, Bed21]. extendable [RRS20]. Extending [BC21]. extensions [ENRV23]. extremal [CLWX25, GHKY20].

facility [CHTW21, LL24a, Man25]. factor [BIM21, GW21]. factorial [ABT21]. factoring [HS25]. factorisations [LMT25]. factorization [HBL24]. failure [GRZ24]. Fair [GJ23, Suk25]. families [Sev20]. family [Sin23]. Fast [MSYY24, FKMS20, KN20]. Faster [Bae22, Tsu20a, Tsu21b, Tsu23a, Tsu23b, Zsc22]. fat [CEP25]. fat-shattering [CEP25]. Fault [SM21]. Fault-tolerant [SM21]. faulty [DS21, GKL^{+}23, LL25, SM21, WY20]. February [Ano20n, Ano21l, Ano23f, Ano24f, Ano25c]. Feistel [SI22a]. few [FGIK24]. field [ID23]. FIFO [Zei23]. filtered [GKP22]. final [Vol23]. finder [CIM20]. Finding [GPWM24, Ish21, Mol22]. fine [Ami21]. fine-grained [Ami21]. finite [CST23, Doy25, GLW23, HBL24, ID23, PS20]. finite-field [ID23]. first [An22, DKS24]. first-order [DKS24]. fixed [Ish21, Ism24]. flip [BCK23b]. flips [DK21]. flow [PZ24]. flowers [DSTZ24]. flows [BS21]. FM [KC21]. FM-index [KC21]. folded [ZC23]. forest [HvIMS25]. forest-based [HvIMS25]. forgeries [DMM20]. forming [PR24]. Formulæ [EIP22]. formulas [Chi20]. formulation [AF20]. forward [LMT25]. Fourier [AY21]. FPT [FGIK24, GJ23, Tsu20b]. fractal [EAE21, MY18]. Fractionally [MS20a]. fractions [PB23]. framework [TP24]. frameworks [ENRV23]. Fréchet [FK25, Mir24]. Free [Kos25, BHMP22, Feg23, Kam21, Mac24, RRS25, Tsu21c]. full [CWW20]. full-round [CWW20]. fully [Ohs21, RT20b, RT23]. function [BFJ22, CWW20]. functions [DKMS24, HW21, HLS20, LPT20, PF23, SI22b, ZKP^{+}24]. Further [WL22].

Gabriel [BMS20]. gadget [Cha25]. game [CHTW21, FPP23, Rab22]. games [DSTZ24, DH24, Goe20, Zim22]. Gap [SPG22, JPV22, JR20, UW21, DGI21]. Gaussian [DFW22, Gia21]. general [CIM20, EK20]. generalization [Rab22]. Generalized [MP23, BS21, KM21, RV23, SI22a, WZ25, ZC23, ZKP^{+}24]. generated [LS23]. generating [CCJS22]. generation [ACCL23]. generators [GKP22, KLM23, Vig20]. generic [HYZ^{+}20]. genetic [CDDN21, CIM20]. Geodesic [AS21]. Geometric [AA22, BBBMS22, DFMHVHT21]. Glauber [JPV22]. global [Pou22, Smy20]. Globally [GM22]. good [DSTZ24]. good-for-games [DSTZ24]. goods [DH24]. GOST [CWW20]. goto [Che23]. GR [CGG^{+}23b]. grained [Ami21]. Graph [CR20, Bod22, CCJS22, DH23, Dra20, Fri21, GH25, HMR24, KM21, OT21, TY23, TV23, ZWWC22]. Graphs [BCD20, AK22a, AFK^{+}24, Ami21, An22, APS25, ACG^{+}24, BMS20, BHMP22, CX21, CLWX25, DFMHVHT21, EA24, Feg23, FHL21, Fri21, GW21, GB21, GGSdS20, HPR22, HPP20, HLS20, HW22, IK22, JJD22, JKL21, KP24, Kno21, LLLW23, MP20, PCC20, PW21, PP24, PF23, RRS20, RR23, Sah22b, Sch24, SS22, Sev20, Sha23, Tan22, Tsu21a, VP20, WHN25, XN20, Zam22, ZWWC22]. gravity [Cha25]. greedy [HK20, WL22, BCKP23]. Grid [LMO^{+}22]. grids [DNS20, Jac21]. group [HBL24, PCC20, Sch21]. groups [HBL24, NS24]. Grundy [VP20]. Guided [BKS23].

half [HKR21, PW21]. half-edges [PW21]. half-integral [HKR21]. Hamilton [DE23]. Hamiltonian [An22, BMS20, DS21, KM21, LL25, Zam22]. Hamiltonian-connectedness [An22]. Hamiltonicity [SM21]. Hanano [Cha25]. handling [YL22]. Hard [MPS22, GPWM24, Man21]. Hardness [BDH21, AC21, BB21, BDK^{+}24, HPP20, LW23, Man24, Zei23]. hashing [MP23]. Hausdorff [FK25]. hazards [KS20]. Heap [EIP22]. Heavy [DE23]. heterogeneous [AM20, Sha23]. heuristics [CCJS22]. hexagonal [Jac21]. hidden [Ism24]. Hierarchical [CR20]. high [AHKBS22, Gia21, YKC24]. high-dimensional [AHKBS22]. higher [BMS20, Mir24, VPT24]. higher-order [BMS20]. Hitting [vBS20]. hole [WYZ^{+}24]. homogeneous [Chi20]. hop [HPP20]. Horton [Bie22]. hub [FRRT22]. Huffman [GK23]. hull [ACG^{+}24]. hunt [PP24]. hypercube [LL25]. hypercubes [DS21, SM21]. hypergraphs [CFHH21, GPWM24, HS24b, Kam23, QW24]. hypernetworks [GPWM24]. hyperplane [AHKBS22]. hyperplanes [AA24]. ideal [Xia20]. identical [MSS24]. identification [GRZ24]. Identity [SP20]. IIoT [PK24, ZXY^{+}22]. im [SI22a]. improve [HS24a]. Improved [ABT21, CGG^{+}23a, Dür23, GKP22, Man25, MS20a, ZCWW21, AK22a, BCK23b, Chi20, CEP25, GHN^{+}25, KLM23, Liu23, LMT25, WXC24, XK22]. improvements [ZXY^{+}22]. Improving [VDPT25, SI22a]. Inapproximability [QW24, DH24]. inclusion [BDK^{+}24]. inclusion-wise [BDK^{+}24]. increasing [Lev22]. Incremental [WY20]. independence [JPV22, RR23]. Independent [LMMZ20, AC21, QW24, Sel24]. index [An22, KC21]. indifference [MP20]. individual [Chi20]. indivisible [Suk25]. induced [WXC24]. Inductive [EIP22, HT21]. Inf [RT23]. infinite [DLN^{+}23, Sha20, Zim22]. infinitely [ZQKL24]. Information [BH22a, Sup22, WL21]. information-theoretic [Sup22]. injective [FHL21]. inner [VPT24]. input [Sak20]. input-driven [Sak20]. Insertion [LM22]. Instability [Gia21]. integer [PH23]. integers [DFW22, WW24]. integral [BS21, HKR21]. integrality [JR20]. interactive [MG20]. interoperability [LLP20]. Interrupt [BH22b, BH22a]. interval [GM24]. intractability [Sin23]. invariants [vIMM23]. isogenous [HJHZ22]. isogeny [HJHZ22]. isogeny-based [HJHZ22]. isolation [CLWX25]. Isomorphism [CFHH21]. items [JZ22, Suk25]. iterativity [PR24].

January [Ano20o, Ano21m, Ano22i,

Ano23g, Ano24g, Ano25d]. Jha [KK21]. job [Lev22, Sin23, YHK24]. job-dependent [YHK24]. Jr [ZCWW21]. July [Ano20p]. June [Ano20q, Ano21n, Ano22j].

Keccak [ZCWW21]. Keccak-MAC [ZCWW21]. Kernel [Tsu21c, BCK23b, KLM23]. kernels [BCK^{+}23a, CGG^{+}23a, Tsu24, vBS20]. Ketje [ZCWW21]. Ketje-Jr [ZCWW21]. key [TP24, ZXZ^{+}23]. Kinetic [Aba21]. knapsack [ABM20]. knowledge [MG20]. König [RRS20]. Kruskal [BRS21].

labeled [BCEM24]. labeling [DLN^{+}23]. Labelled [HLS20]. labelling [TZZ25]. labelling-based [TZZ25]. lambda [Mac24]. lambda-calculus [Mac24]. language [MS24b]. languages [LS23, Sak20]. Laplacian [GB21]. large [Sha23, WXC24]. larger [BC21]. lateness [WTP24]. lattice [LTT23, PH23, ZY23]. lattice-based [LTT23]. lattices [AC21, SP20, WL21]. layer [Ism24]. Lazy [KN20]. LBFS [Sch24]. LD [RT20a]. leakage [ZXZ^{+}23]. learnability [tCFJL24]. learnable [KP21]. Learning [LdOOW24, PZ24, YHK24]. Learning-augmented [PZ24]. leasing [ZXH20]. least [BS23]. left [ABT21]. lemma [AK22b]. lemmas [Jai20]. length [LMT25, LF20, YKC24]. length- [LF20]. Lengths [PS20, GHKY20, Tan22]. Lett [RT23]. Letters [BH22a]. level [FRRT22, vIKMN22]. level-2 [vIKMN22]. leveling [DGI21]. Leveraging [DGM25]. lexicographic [GS22]. Lie [HBL24]. light [DE23]. like [ZCWW21]. line [BDBK^{+}25, LMO^{+}22, MS24a, Sha20, WYZ^{+}24]. Linear [DKMS24, DFL^{+}20, Man21, APEiC22, APEiC24, Dra20, EA24, EK20, HS20, JA20, KLM23, Kno21, LZG22, MT23, MSYY24, QW24, vBS20]. linear-time [EA24, MT23]. linearizations [APEiC22]. lines [JR23]. List [BHMP22]. listing [Liu23]. lists [DKMS24]. load [HW22]. loading [Klu24]. Local [Kno21, MSYY24, Pou22]. location [CHTW21, LL24a, Man25]. log [DKP^{+}20]. logarithm [WZ25]. logarithmic [DKP^{+}20]. logarithms [WZDZ22]. Logic [EIP22, Eng21, GM24, MT21, MW23, EP23]. logics [Zim25]. Long [JWZ24, Sha23]. Longest [BIM21, BIK23, KHO21, Sha20, Bli20, DBRB21]. lookahead [Zim22]. loss [ACG23]. lost [CT20]. Lot [MMS20]. Lovász [AK22b]. low [PH23]. Lower [GH25, BKK23, CHTW21, Chi20, EHL^{+}21, GHN^{+}25, GL25, LXZW23, Man25, Sah22a, Sup22, YQL22, ZY23]. LP [JR20]. LPT [MSS24]. LTL [BFM23]. Luby [GTZ25]. Lyubashevsky [TP24].

Ma [EAE21]. MAC [ZCWW21]. machine [AF20, GLW23, LG23, LZG22, MMS20, Sin23, WTP24, YHK24]. machines [AM20, Lev22]. macro [MS23]. made [CM22, GTZ25]. main [EAE21]. majority [BS23]. makespan [AM20, LG23]. makespans [LZG22]. manipulating [MK20]. Mansour [SI22b]. many [FGIK24, IK22, Zam22]. mapping [BCEM24]. maps [KS23]. March [Ano20r, Ano22k, Ano23h, Ano24h]. Markov [FHL^{+}23, Vah24]. masked [ID23]. matching [Feg23, Kam21, KN20, KC21, KKNS23, MMCH20, MSHS23, WHN25, Zsc22]. matching-cut [Feg23]. matchings [AABC20, BMS20, PW21, TV23]. matrices [AY21, APZT22, Pud22]. Matrix [SPG22, KS22, VDPT25, WW24, WZ25]. Max [Sta22]. Max-CSPs [Sta22]. Maximal [MW23]. maximization [ABM20, BCKP23]. Maximum [vdHKL^{+}20, APEiC24, HR20, LG23, LL24a, MSHS23, PZ24, WTP24, RSRM23]. Maximum-area [vdHKL^{+}20]. may [KN20, Ano20s]. MDS [WZ25]. means [GOR^{+}22, Pou22]. means/median [BFJ22]. mechanisms [CHTW21]. median [BFJ22]. meet [Sha21, TZZ25]. memory [DGI21, RT20b, RT23]. Mesnager [LPT20]. Mesosome [CSS23]. method [JA20, VDPT25, WZ25]. metrics [HHMM20]. Mim [BHMP22]. Mim-width [BHMP22]. Min [Dür23, Sta22]. Min-CSPs [Sta22]. Min-Plus [Dür23]. Minimal [JKL21, BDK^{+}24, GM22, Klu24, MW23, Zim22]. minimize [LZG22, MMS20]. minimizing [FKMS20, LG23, WTP24]. Minimum [APEiC22, AK22a, HKR21, LMO^{+}22, Mas21, Nut25, Sha23, TV23]. minor [DHW22]. MinSat [Fio22]. MIS [Ami21, GTZ25]. mismatches [Sok20]. mixed [MS24a]. modalities [GM24]. model [CIM20, MT21]. models [EAE21, MY18, WL23]. modification [Tsu20a]. Modified [GHKY20]. Modular [AR22, Mac24]. monadic [Eng21]. Monochromatic [AABC20, JR23]. monoids [Kos25]. monomial [ZKP^{+}24]. monotone [BCKP23, Dür23, Mas21]. monotonic [PB23]. Morse [Bli20]. most [TY23]. MST [BRS21]. mule [YL22]. multi [FGS23, Ito25, RT20a]. multi-branching [Ito25]. multi-commodities [FGS23]. multi-processor [RT20a]. multiplayer [Goe20]. multiple [KS23, WZ25]. multiplication [VDPT25]. multiplier [ID23]. multiprocessor [FKMS20]. multisignature [LTT23]. multivariate [OPD23]. Multiway [SPG22]. mutated [HR20]. Mutual [RT20b, RT23].

nail [Vol23]. naive [ENRV23]. Nash [DH24]. near [Mir24]. near-neighbors [Mir24]. nearest [Gia21]. nearly [BKK23]. needed [Zim22]. negative [LdOOW24]. neighbor [Gia21, TF23]. neighborhood [LF20]. neighborhood-constrained [LF20]. neighbors [Mir24]. nesting [MS23]. Network [GRZ24, MS20a, FPP23, GB21, HvIMS25, Sah22b]. network-based [GB21]. networks [BCV21, Ism24, LLC21, MMCH20, Sel24, Ste20, YQL22, Zei23, vIKMN22]. neural [Ism24]. nilpotency [NS24]. no [Ray24]. node [GRZ24]. Noisy [Alw20]. Non [EIP22, GK22, MM20, WZDZ22, Fio22, Lev22, MSS24, MG20, PK23, Sha20, Zei23, tCFJL24]. non-clausal [Fio22]. non-efficient [tCFJL24]. Non-Established [EIP22]. non-FIFO [Zei23]. non-identical [MSS24]. non-increasing [Lev22]. non-interactive [MG20]. Non-Preemptive [MM20]. non-primitive [PK23]. Non-uniform [GK22, WZDZ22, Sha20]. nonces [LL24b]. noncrossing [OT21]. nondecreasing [KL20]. nondeterministic [PS20]. nonlinear [HW21, ZKP^{+}24]. norms [KS22]. note [AC21, Ami21, BS21, Bod22, Bra22, CDP23, Feg23, FPP23, Fuj23, IS22, Jai20, JR20, LT25, Liu23, LS23, MMCH20, Man24, MK20, PH23, YHK24]. November [Ano20t, Ano22l]. NP [GPWM24, UW21, Zei23, ZQKL24]. NP-complete [UW21, ZQKL24]. NP-hard [GPWM24]. NP-hardness [Zei23]. NTRU [SP20, ZY23]. number [ACG^{+}24, Bie22, DFMHVHT21, EAE21, GW21, GKP22, KK21, Kno21, LMO^{+}22, MY18, MSS24, MMS20, Ohs21, Pou22, Ruk20, Sah22a, Sev20, Sup22, TY23, Vig20]. numbers [Jac21, RR23]. numerosity [Cha25].

obfuscator [Vol23]. objects [AA22, AA24]. oblivious [GIR20]. OCB3 [LL24b]. October [Ano20u, Ano21o]. odd [TY23]. offline [BEL20]. On-line [WYZ^{+}24]. one [HKP21, Liu23]. one-counter [HKP21]. Online [AHKBS22, Fri21, LZG22, ZXH20, BEL20, DGI21, ZZLC22]. open [LPT20, TGS25]. operations [XN20, ZZ21]. operator [Moo22]. opinion [PR24]. opportunity [ZXH20]. opposite [CHTW21]. Optimal [CM22, GKL^{+}23, Ito25, vBS20, BCV21, CST22, CCJS22, HKR21, LTT23, LP22, Sup22, ZXZ^{+}23]. Optimal-size [vBS20]. optimality [UNSI24]. optimization [Kam23]. Optimizing [HJHZ22]. Oracle [MG20].

ORANGE [DMM20]. Order [KKNS23, BMS20, DKS24, Eng21, VPT24]. Order-preserving [KKNS23]. ordered [BRS21]. orderings [HT21]. orders [GS22, MMS20]. Ore [WQ20]. Ore-type [WQ20]. orientability [KP24]. oriented [ID23]. origin [MS23]. orthogonal [AA22, AY21, Tsu20b]. other [AA24]. Overlap [CR20, Vig20]. overlapping [BIM21].

PAC [tCFJL24]. Packing [JZ22, JZ23, HLS20, Ray24]. packings [GK22]. pairs [LP22]. pairwise [HPR22, Kam21, XN20]. palindrome [GS21]. palindromes [Sok20]. Palindromic [MWN^{+}22]. paradigm [CDDN21]. paradox [DFLS23]. parallel [AS21, AA24, CL23, ID23, LG23]. parallel-batch [LG23]. parameterized [DGM25, KC21, Ohs21, RSRM23, Tsu21b, XK22, Zsc22]. parameters [GKNS23]. Pareto [BCV21]. Parikh [APZT22]. parity [GKP22]. partial [BS21]. partially [BRS21]. partially-ordered [BRS21]. partition [BCD20, CDP23, PCC20]. partitioning [GGSdS20, JR23]. partitions [ACCL23]. path [BKK23, DLN^{+}23, LMO^{+}22, LF20, Zei23, Tsu23b]. Paths [JWZ24, BCK^{+}23a, BCV21, CGG^{+}23a, DE23, DS21, KL20, Nut25, Sha23, WY20]. pathways [HR20]. pathwidth [Bie22]. pattern [KKNS23]. Paxos [Sut20]. Penalty [SPG22]. perfect [Goe20, TV23, ZKP^{+}24]. Periodic [MM20, BIK23, PB23]. permutations [ÁRCLM^{+}22, BCKV21, UNSI24]. perspective [BHMP22]. phylogenetic [vIKMN22]. piccolo [UNSI24]. piccolo-type [UNSI24]. planar [Bae22, DKP^{+}20, HMR24]. planarity [APEiC24, UW21]. plane [AABC20, ÁRCLM^{+}22, DK21]. Plus [Dür23]. point [AABC20, ÁRCLM^{+}22, Bae22, Ish21, OT21]. points [AS21, JR23]. polygon [MRD25, vdHKL^{+}20]. polygonization [CDDN21]. polygons [Bae22, BCK23b]. polylogarithmic [BN22]. Polynomial [BCK^{+}23a, FGS23, MM20, vIMM23, BMWW22, BH22a, MS24b, Ohs21, OFA21, TV23, YL22, BH22b]. population [CL23]. Poset [OFA21]. positive [LdOOW24]. positivity [Vah24]. possibility [SI22a]. potentially [Pud22]. power [DLN^{+}23, LM22, Nut25]. power-weight [LM22]. powers [DG23]. powers-of-two [DG23]. PPS [HHT23]. Practical [DMM20]. Pradhan [KK21]. preclusion [MMCH20]. preconditioned [Che23]. Preemptive [AM20, MM20, Lev22, Sin23]. preferences [Kam21]. preferred [ENRV23]. prefix [AR22, LMT25, LP22]. prefix-constrained [AR22]. prefix-stable [LMT25]. prefixes [GS21]. Prenex [DKS24]. presence [GKL^{+}23]. preserving [Bod22, KKNS23]. previous [BIM21]. primal [MSYY24]. primal-dual [MSYY24]. primitive [PK23]. priori [FS20]. prisms [KM21]. privacy [GKNS23]. private [Man25]. privileged [Ruk20]. Probabilistic [CT20, WL23, Bha22, GKNS23, Zim25]. probabilities [DG23]. probability [Vig20]. Problem [SPG22, AA24, APEiC24, AF20, BKK23, BDH21, BCK23b, CDP23, DHW22, DBRB21, EHL^{+}21, EA24, FS20, FKMS20, GLW23, HR20, HBL24, Kam21, Klu24, Kos25, LPT20, LL24a, LT25, MMCH20, MRD25, MSYY24, PRM24, PH23, Sch21, WZDZ22, WZ25, vBS20, OFA21, RSRM23, ZZLC22]. problems [BRS21, HKP21, HMR24, HS24b, Kam23, Lab24, LF20, MS24a, MPS22, Tsu20a, Tsu21a, Tsu24, Zei23]. Process [RT23, Sha20]. processes [Bha22, FHL^{+}23, WYL24]. Processing [BH22a, WTP24]. processor [RT20a]. processors [MSS24]. product [Hua23, VPT24, Dür23]. products [Kos25, WY20, WW24]. program [CIM20]. programming [CIM20, JS20]. projective [APEiC22]. projectivity [APEiC24]. proof [AK22b, BB21, CHTW21, EAE21, LdOOW24, MT21, MS20b]. proper [HvIMS25]. properties [DKS24]. property [MT21]. proportionally [BCD20]. protocol [GIR20]. protocols [CL23]. proxy [PK24, ZXY^{+}22]. prune [KLM23]. pseudorandom [SI22b, Vig20]. Public [ZXZ^{+}23, DHP^{+}22, DH24, LKC22]. public-coin [DHP^{+}22]. Public-key [ZXZ^{+}23]. pure [JS20]. purpose [CIM20]. pushdown [EHL^{+}21, HKP21]. Puzzle [Cha25].

QBF [BB21]. quadratic [GHKY20]. Quantitative [FHL^{+}23]. Quantum [SI22b, MG20, WYL24]. quaternary [YKC24]. queries [DKP^{+}20, tCFJL24]. query [PH23]. questions [Kos23]. queueing [GB21]. quickly [MPS22]. QUIXO [MT20]. quorum [Sah22a].

R [CGG^{+}23b]. R3 [AS21]. RAC [RE20]. radio [DLN^{+}23]. Raiders [CT20]. railways [JN25]. ramp [EK20]. Ramsey [LMT25]. Random [BCKP23, CFHH21, GKP22, MS20b, Sha23, Vig20]. Randomized [BEL20, HHT22, DHP^{+}22, FPP23, Ito25]. range [DKP^{+}20]. Rank [ZZ21]. Rankin [LXZW23]. Ranking [Gab24, MT23]. rate [PB23]. ratio [RT20a]. rational [Kos23]. reachability [BH22a, BH22b, FQSW20, HKP21]. reachable [Ohs21]. real [MS20b, PB23, TV23]. real-time [PB23]. real-valued [MS20b]. realizing [HHMM20]. reasoning [EP23]. reciprocal [An22]. recognition [FGS23]. Recognizing [AFK^{+}24, HHMM20, Sch24, RR23]. reconstructing [vIKMN22]. Reconstruction [OT21, MMHX20]. recovery [TP24]. rectangle [GL25]. rectangular [Dür23, WYZ^{+}24]. Recursion [Mac24]. Recursion-free [Mac24]. recursive [Man24, Ste20]. Red [AA24]. Reduced [TZZ25, ZCWW21]. reduced-round [ZCWW21]. Reducing [HS25]. Reduction [JWZ24, WHN25, XN20]. Refined [WZDZ22, GOR^{+}22]. registers [HHT22]. regraft [KLM23]. Regular [BY24, LMT25, Ami21, HHT22, LS23, MS24b, Sak20, TGS25]. regularity [Sta22]. Relating [HMR24]. relation [FQSW20, Jai20]. relative [WL22]. release [Sin23]. ReLU [Sel24]. remainder [LKC22]. Remark [HS20]. rendezvous [PP24]. Representations [WL21, Bie20]. representatives [BBBMS22]. Representing [ÁRCLM^{+}22]. requirement [CDP23]. Residual [BCKP23]. residuation [GS22]. residue [GHKY20]. residues [ABT21]. resilience [ZXZ^{+}23]. resolution [Bie20, BY24]. Resource [HW21]. restricted [BCKV21, CST23, JR20, PW21]. restriction [LT25]. result [EP23, EAE21]. Results [Sch21, Gia21, HPR22, HPP20, Liu23, PF23, WL22]. revealing [AK22b]. Reverse [HK20]. reversing [HS24a]. review [Suk25]. Revisited [Ray24, GIR20, GK23, MP23, Tsu22, WZDZ22, vdHKL^{+}20]. Revisiting [BH22a, BH22b, ZY23]. rewrite [AR22]. ring [GKP22, Klu24, OPD23, WW24]. rings [Ste20]. Robbins [Sha21]. robots [GKL^{+}23]. Robust [Lev22, WTP24, Zim25]. role [Sch21]. Roman [PF23]. rooted [Bie22, HW21, KLM23, Mas21]. Rotating [Rab22]. round [CWW20, LTT23, Liu23, SI22a, ZCWW21]. rounding [HKR21]. routing [FS20, SS22]. RSA [LKC22]. Rules [EIP22]. Runtime [Doe21].

safety [AT24, DKS24]. Salesman [ZZLC22]. sample [HHT23, Sel24]. sampling [DFW22, HHT23]. Santa [JR20]. satisfactory [CDP23]. satisfiability [HMR24]. satisfying [JKL21]. scaling [FKMS20, KKNS23]. Schatten [KS22]. Schedulability [MM20, PB23]. scheduling [AM20, AF20, Lev22, LG23, LZG22, MSS24, MMS20, RT20a, Sin23, WTP24, YHK24, YL22]. scheme [JA20, LTT23, PK24, SP20, ZXZ^{+}23, ZXY^{+}22]. schemes [EK20, OPD23]. Schröder [BCKV21]. Schulze [MK20]. scissors [Tsu20b]. search [AHKBS22, DBRB21, GKL^{+}23, Gia21, Mir24]. searches [ZWWC22]. second [Eng21, WL22]. second-order [Eng21]. secret [EK20, JA20, MMHX20, PCO20]. secure [HYZ^{+}20, KK21, MMHX20, PRM24, Vol23]. secured [LKC22]. Security [JA20, PK24, CK23, EK20]. Segment [Bie20]. segments [BDBK^{+}25, MS24a]. select [ZZ21]. Selective [Jai20]. self [EAE21, GTZ25, GHKY20, GKP22, MY18, MS23]. self-dual [GHKY20]. self-nesting [MS23]. self-similar [EAE21, MY18]. self-stabilizing [GTZ25]. self-timed [GKP22]. Semantic [DSTZ24]. semantics [TZZ25, WYL24]. semi [KP24, WL21]. semi-lattices [WL21]. semi-transitive [KP24]. semicomplete [Xia20]. semigroups [Kos25]. semirings [BRS21]. separability [AA22]. Separation [EIP22, GM24, MSHS23, EP23]. separations [MG20]. separators [BDK^{+}24, Mol22]. September [Ano20v, Ano21p]. Sequence [SPG22]. sequences [BC21, YKC24]. serial [WTP24]. serial-batch [WTP24]. series [Kos23]. serving [ACG23]. Set [AA24, vBS20, AABC20, AK22a, Bae22, BN22, CST23, Fuj23, Kno21, OT21, PRM24, Tsu23b]. Sethi [RT20a]. sets [ÁRCLM^{+}22, ENRV23, GGSdS20, QW24, ZQKL24]. setup [Sin23]. several [Sev20]. shalt [TF23]. shared [RT20b, RT23]. sharing [EK20, JA20, MMHX20, PCO20]. sharp [Sah22a]. sharper [LXZW23]. shattering [CEP25]. short [LL24b, YHK24]. shortest [AC21, BCK^{+}23a, WY20, Zei23]. side [ID23]. side-channel [ID23]. sieve [HS25]. signature [OPD23, PK24, SP20, TP24, ZXY^{+}22]. Signed [DNS20, Jac21]. similar [EAE21, MY18]. simple [ABM20, BB21, BCK23b, KL20, Lou20, MT21, PW21, XK22]. Simpler [KC21]. simplest [GIR20]. Simplicity [AT24]. Simplified [DG23, DGM25]. simulates [BY24]. simultaneous [DK21]. single [AF20, Ism24, JZ22, JZ23, LZG22, MMS20, Sin23, YHK24]. single-machine [AF20, LZG22, Sin23, YHK24]. size [CDP23, HHT23, Sel24, vBS20]. size-independent [Sel24]. sizes [Lev22]. Sketch [VPD24, VPT24]. Skolem [Vah24]. skyline [DKP^{+}20]. sliding [KS22, MWN^{+}22]. sliding-window [KS22]. Slightly [Chi20]. Small [KM21, Bie20, MT21, MSS24]. Smaller [Tsu24]. Smoothness [KS22]. Socially [GJ23]. solutions [CST22, Klu24]. solve [BRS21]. solver [GB21]. solving [CST23]. Some [XN20, APS25, Kos23, Lab24, VP20]. sort [KM21]. sortable [BCKV21]. Sorting [JS20]. Space [GK23, AHKBS22, HS25, KHO21, LP22, vBS20]. Space-efficient [GK23]. span [YL22]. Spanners [FK25, AS21, CCJS22]. spanning [Dra20, EAE21, HKR21, LLC21, MY18, Mas21, OT21]. sparse [LLLW23, WHN25]. sparsification [Bod22]. Sparsifying [VPD24]. Spectral [JPV22, GHN^{+}25]. speed [FKMS20, JS20]. speed-scaling [FKMS20]. split [KP24, LLC21, MP20, Tsu21a]. split-indifference [MP20]. split-star [LLC21]. Splitting [ZQKL24]. square [GS21]. squares [PK23]. stabilizing [GTZ25]. stable [BS23, LMT25, WHN25]. stably [WL21]. stacks [BCKV21]. star [LLC21]. Start [DGI21]. Start-Gap [DGI21]. stash [MP23]. Static [Moo22]. Statistical [Bed21]. statistically [Vol23]. statistically-secure [Vol23]. Steiner [FRRT22, HS24b, ZZLC22]. Strahler [Bie22]. straight [LMO^{+}22]. straight-line [LMO^{+}22]. Strategies [CCJS22]. strategy [CHTW21, ZXH20]. strategy-proof [CHTW21]. streams [KS22]. string [KC21]. Strong [EK20, LW23, LLLW23, MMCH20]. strongly [HKR21, WHN25]. structures [EK20, SI22a]. Subadditive [MS20a]. subclasses [VP20]. subcodes [HS20]. subcubes [SM21]. subcubic [PW21, RE20]. subexponential [FGIK24]. subgame [Goe20]. Subgraph [BDH21, APS25, Bra22, HKR21]. subgraphs [BCD20, DH23, Sha23, WXC24, Zam22]. subject [ABM20, HW21]. sublinear [KHO21]. submodular [ABM20, BCKP23, PH23]. subsequence [Bli20, DBRB21, KHO21]. subsequences [BIK23, Vig20]. subset [AF20]. subset-sum [AF20]. subspace [JA20]. subtree [KLM23, Pou22]. subtrees [Pou22]. suffix [LP22]. suffix/prefix [LP22]. Sum [SI22b, AF20, HW21]. Super [LLC21, ZC23]. superimposed [RV23]. supports [Kos23]. Surveying [Smy20, SC22]. swap [DFL^{+}20, ZC23]. Symbolic [EIP22]. symmetric [ABM20]. symmetry [Doe21]. Synchronizing [BFM23]. synthesis [FHL^{+}23]. synthesizer [CIM20]. System [MM20, CST23, Sup22]. systems [AR22, BBBMS22, HKP21, LS23, RT20b, RT23, WL21].

Table [Rab22]. tableau [Fio22]. Tai [BCEM24]. tardy [MMS20]. Task [MM20]. tasks [PB23]. teaching [Man24]. teams [TGS25]. techniques [Bra22]. temporal [GM24, Mol22, Zim25, Zsc22]. term [AR22]. termination [AR22]. terms [Sup22]. terrains [KS23]. Test [MM20, Bha22]. Testing [UW21]. their [VP20]. theorem [DFLS23, LKC22, Sax21]. theoretic [Sup22]. theory [EP23]. There [Ray24]. Thou [TF23]. three [Sax21]. Threshold [MMHX20, DKMS24, JA20, OPD23]. Thue [Bli20]. thy [TF23]. Tight [CHTW21, DH24, GW21, GJ23, HKP21, Sup22, BKK23, ZZLC22]. Tighter [WZ25]. tiling [GL25]. Time [MM20, APEiC22, BCV21, CL23, DFL^{+}20, DKP^{+}20, Dra20, DFW22, EA24, FKMS20, GM24, LP22, MS24b, MT23, Mol22, OFA21, PB23, YL22, Zei23, vBS20]. time-dependent [BCV21, Zei23]. Timed [BH22b, BH22a, FQSW20, GKP22]. times [Sin23, Zei23]. tolerant [SM21]. Tomography [GRZ24]. tool [LMT25]. tools [Sup22]. Top [Doy25, MS24b]. Top-down [Doy25, MS24b]. toroidal [TY23]. tracking [BCK^{+}23a, CGG^{+}23a]. tractability [BDH21]. Tradeoff [AM20]. transducers [MS23]. transfer [GIR20]. transform [AY21]. Transforming [DK21]. transitive [KP24]. Transposition [SPG22]. transversals [LW23]. travel [Zei23]. Traveling [ZZLC22]. TrCBC [CK23]. treasure [PP24]. tree [BMWW22, CCJS22, Dra20, LW23, MS23, MS24b, Mas21, Pud22]. trees [APEiC22, APEiC24, Bie22, BS21, BCEM24, DKMS24, Doy25, EAE21, Fri21, FRRT22, HW21, Ito25, LMO^{+}22, LM22, MY18, MWN^{+}22, OT21, Pou22, Sah22a, Sch24]. treewidth [GHN^{+}25, SS22]. triangle [vdHKL^{+}20]. triangular [Jac21]. triangulations [DK21]. trinets [vIKMN22]. tuple [JJD22]. twin [Kno21]. Two [FRRT22, AF20, BCKV21, BCD20, CHTW21, DG23, LZG22, MT21, MS24a, Tsu24, WQ20]. two-agent [AF20]. two-center [MS24a]. Two-level [FRRT22]. two-opposite-facility [CHTW21]. two-variable [MT21]. type [AF20, HBL24, Moo22, OT21, UNSI24, WQ20]. types [GLW23].

unambiguous [IK22]. Unbiased [VPT24, BBBMS22]. unbordered [Gab24]. unbounded [LG23]. undecidability [EP23]. Undecidable [EIP22]. unicyclic [Fri21]. uniform [CFHH21, CEP25, FS20, GW21, GK22, Lev22, Sch21, Sha20, WZDZ22]. Uniqueness [AT24]. unit [FS20]. universal [DKS24]. unordered [BCEM24]. unpredictability [PCO20]. unranking [Gab24, MT23]. unrefinable [ACCL23]. unsplittable [JN25]. update [MSYY24]. Upper [Ruk20, GH25, YQL22]. upward [Bie22]. used [HS25, LL24b]. useful [Pud22]. using [AA22, JA20, LP22, LKC22, PB23, SPG22, Sup22, VDPT25].

validity [BMWW22]. valued [MS20b]. variable [MT21]. variables [CST23, MS20b]. variate [VDPT25]. Variational [Ste20]. VAS [EHL^{+}21]. Vector [MP20, AC21, Ray24]. vehicle [FS20]. verifiability [Smy20, SC22]. Verification [ACCL23]. versus [LMMZ20]. Vertex [GRZ24, Tsu23a, Zam22, DHW22, HT21, RRS20, Tsu24]. Vertex-connectivity [GRZ24]. vertex-deleted [Zam22]. vertex-minor [DHW22]. vertices [Ohs21, ZWWC22]. very [MPS22, WHN25]. via [Doe21, PW21, VPT24, WY20, WZ25]. viewpoints [KS23]. virtual [GLW23]. visibility [KS23]. VNP [IS22]. VNP-completeness [IS22]. Voronoi [KS23]. voting [MK20]. vulnerable [FGS23].

waiting [Mol22]. Waypoint [SS22]. weak [Goe20, WL23]. Weakest [Che23]. weakly [Ito25, MS23]. weakly-balanced [Ito25]. weakness [LL24b, PK24]. wear [DGI21]. weight [LM22, MSYY24, TV23, WL22]. Weighted [KP21, Suk25, Tan22, AM20, BCV21, EA24, HKR21, LMMZ20, MMS20, MK20]. weights [Ism24]. well [AFK^{+}24, PCC20, Tan22]. well-covered [Tan22]. well-dominated [AFK^{+}24]. well-equalized [PCC20]. whether [MS24b]. which [CX21]. width [BHMP22]. win [Zim22]. window [KS22, MWN^{+}22]. wise [BDK^{+}24, UNSI24]. without [BCD20, Moo22, Tan22]. WOM [BKS23]. Word [NS24, Kos25, ZZ21]. words [BFM23, Bli20, FRS20, Gab24, PK23, PS20, Ruk20]. Worst [MSS24]. Worst-case [MSS24].

XOR [Jai20].

Yao [EAE21].

Zagreb [An22]. zero [DHP^{+}22, MG20]. zero-error [DHP^{+}22]. zero-knowledge [MG20]. Zone [Sax21].

References

Abidha:2022:GSU

[AA22] V. P. Abidha and Pradeesha Ashok. Geometric separability using orthogonal objects. Information Processing Letters, 176(??):??, June 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                                      /www.sciencedirect.com/
                                                      science/article/pii/S0020019022000023.

Abidha:2024:RBS

[AA24] V. P. Abidha and Pradeesha Ashok. Red Blue Set Cover problem on axis-parallel hyperplanes and other objects. Information Processing Letters, 186(??):??, August 2024. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000152.

Abu-Affash:2020:MPM

[AABC20] A. Karim Abu-Affash, Sujoy Bhore, and Paz Carmi. Monochromatic plane matchings in bicolored point set. Information Processing Letters, 153(??):??, January 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301437.

Abam:2021:KCD

[Aba21] Mohammad Ali Abam. Kinetic collision detection for balls. Information Processing Letters, 171(??):??, October 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902100051X.

Amanatidis:2020:SDA

[ABM20] Georgios Amanatidis, Georgios Birmpas, and Evangelos Markakis. A simple deterministic algorithm for symmetric submodular maximization subject to a knapsack constraint. Information Processing Letters, 163(??):??, November 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300971.

Andrejic:2021:IAL

[ABT21] Vladica Andrejić, Alin Bostan, and Milos Tatarevic. Improved algorithms for left factorial residues. Information Processing Letters, 167(??):??, April 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020301654.

Aggarwal:2021:NCH

[AC21] Divesh Aggarwal and Eldon Chung. A note on the concrete hardness of the shortest independent vector in lattices. Information Processing Letters, 167(??):??, April 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020301526.

Aragona:2023:VGU

[ACCL23] Riccardo Aragona, Lorenzo Campioni, Roberto Civino, and Massimo Lauria. Verification and generation of unrefinable partitions. Information Processing Letters, 181(??):??, March 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000042.

Azar:2023:LSD

[ACG23] Yossi Azar, Ilan Reuven Cohen, and Iftah Gamzu. The loss of serving in the dark. Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000916.

Araujo:2024:HNC

[ACG^{+}24] Julio Araujo, Victor Campos, Darlan Girão, João Nogueira, António Salgueiro, and Ana Silva. On the hull number on cycle convexity of graphs. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000637.

Avolio:2020:SST

[AF20] Matteo Avolio and Antonio Fuduli. A subsetsum type formulation of a two-agent single-machine scheduling problem. Information Processing Letters, 155(??):??, March 2020.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019301693.

Agrawal:2024:RWD

[AFK^{+}24] Akanksha Agrawal, Henning Fernau, Philipp Kindermann, Kevin Mann, and Uéverton S. Souza. Recognizing well-dominated graphs is coNP-complete. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000625.

Antoniadis:2022:OSH

[AHKBS22] Antonios Antoniadis, Ruben Hoeksma, Sándor Kisfaludi-Bak, and Kevin Schewior. Online search for a hyperplane in high-dimensional Euclidean space. Information Processing Letters, 177(??):??, August 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000199.

Abu-Khzam:2022:IEA

[AK22a] Faisal N. Abu-Khzam. An improved exact algorithm for minimum dominating set in chordal graphs. Information Processing Letters, 174(??):??, March 2022.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001216.

An:2022:DRP

[AK22b] Hyung-Chan An and Robert Kleinberg. A diameterrevealing proof of the Bondy– Lovász lemma. Information Processing Letters, 174 (??):??, March 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001095.

Alweiss:2020:NCD

[Alw20] Ryan Alweiss. Noisy corruption detection. Information Processing Letters, 155(??):??, March 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301802.

Aggarwal:2020:PSA

[AM20] Vaneet Aggarwal and Ruijiu Mao. Preemptive scheduling for approximate computing on heterogeneous machines: Tradeoff between weighted accuracy and makespan. Information Processing Letters, 153(??):??, January 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901930153X.

Amiri:2021:NFG

[Ami21] Saeed Akhoondian Amiri. A note on the fine-grained complexity of MIS on regular graphs. Information Processing Letters, 170 (??):??, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000375.

An:2022:FZI

[An22] Mingqiang An. The first Zagreb index, reciprocal degree distance and Hamiltonianconnectedness of graphs. Information Processing Letters, 176(??):??, June 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000047.

Anonymous:2020:A

[Ano20a] Anonymous. April 2020. Information Processing Letters, 156(??):??, April 2020. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2020:D

[Ano20b] Anonymous. December 2020. Information Processing Letters, 164(??):??, December 2020. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2020:EBa

[Ano20c]  Anonymous.  Editorial

Board. Information Processing Letters, 153(??):??, January 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301644.

Anonymous:2020:EBb

[Ano20d]  Anonymous.  Editorial

Board. Information Processing Letters, 154(??):??, February 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301759.

Anonymous:2020:EBc

[Ano20e]  Anonymous.  Editorial

Board. Information Processing Letters, 155(??):??, March 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301917.

Anonymous:2020:EBd

[Ano20f]  Anonymous.  Editorial

Board. Information Processing Letters, 156(??):??, April 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300090.

Anonymous:2020:EBe

[Ano20g]  Anonymous.  Editorial

Board. Information Processing Letters, 157(??):??, May 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300193.

Anonymous:2020:EBf

[Ano20h]  Anonymous.  Editorial

Board. Information Processing Letters, 158(??):??, June 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300387.

Anonymous:2020:EBg

[Ano20i]  Anonymous.  Editorial

Board. Information Processing Letters, 159–160(??): ??, July 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300569.

Anonymous:2020:EBh

[Ano20j]  Anonymous.  Editorial

Board. Information Processing Letters, 161(??):??, September 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300739.

Anonymous:2020:EBi

[Ano20k]  Anonymous.  Editorial

Board. Information Processing Letters, 162(??):??, October 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300910.

Anonymous:2020:EBj

[Ano20l]  Anonymous.  Editorial

Board. Information Processing Letters, 163(??):??, November 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301022.

Anonymous:2020:EBk

[Ano20m]  Anonymous.  Editorial

Board. Information Processing Letters, 164(??):??, December 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301198.

Anonymous:2020:F

[Ano20n] Anonymous. February 2020.

Information Processing Letters, 154(??):??, February 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2020:Ja

[Ano20o] Anonymous. January 2020. Information Processing Letters, 153(??):??, January 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2020:Jc

[Ano20p] Anonymous. July 2020. Information Processing Letters, 159–160(??):??, July 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2020:Jb

[Ano20q] Anonymous. June 2020. Information Processing Letters, 158(??):??, June 2020. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2020:Ma

[Ano20r] Anonymous. March 2020. Information Processing Letters, 155(??):??, March 2020. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2020:Mb

[Ano20s] Anonymous. May 2020. Information Processing Letters, 157(??):??, May 2020. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2020:N

[Ano20t] Anonymous. November 2020.

Information Processing Letters, 163(??):??, November 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2020:O

[Ano20u] Anonymous. October 2020.

Information Processing Letters, 162(??):??, October 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2020:S

[Ano20v]  Anonymous.  September

2020. Information Processing Letters, 161(??):??, September 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2021:Aa

[Ano21a] Anonymous. April 2021. Information Processing Letters, 167(??):??, April 2021. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2021:Ab

[Ano21b] Anonymous. August 2021.

Information Processing Letters, 169(??):??, August 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2021:D

[Ano21c] Anonymous. December 2021. Information Processing Letters, 172(??):??, December 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2021:EBb

[Ano21d]  Anonymous.  Editorial

Board. Information Processing Letters, 165(??):??, January 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301460.

Anonymous:2021:EBc

[Ano21e]  Anonymous.  Editorial

Board. Information Processing Letters, 166(??):??, February 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301587.

Anonymous:2021:EBd

[Ano21f]  Anonymous.  Editorial

Board. Information Processing Letters, 167(??):??, April 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000041.

Anonymous:2021:EBe

[Ano21g]  Anonymous.  Editorial

Board. Information Processing Letters, 168(??):??, June 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000168.

Anonymous:2021:EBf

[Ano21h]  Anonymous.  Editorial

Board. Information Processing Letters, 169(??):??, August 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000429.

Anonymous:2021:EBg

[Ano21i]  Anonymous.  Editorial

Board. Information Processing Letters, 170(??):??, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000570.

Anonymous:2021:EBh

[Ano21j]  Anonymous.  Editorial

Board. Information Processing Letters, 171(??):??, October 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000673.

Anonymous:2021:EBi

[Ano21k]  Anonymous.  Editorial

Board. Information Processing Letters, 172(??):??, December 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021001009.

Anonymous:2021:F

[Ano21l] Anonymous. February 2021. Information Processing Letters, 166(??):??, February 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2021:Jb

[Ano21m] Anonymous. January 2021. Information Processing Letters, 165(??):??, January 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2021:Jc

[Ano21n] Anonymous. June 2021. Information Processing Letters, 168(??):??, June 2021. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2021:O

[Ano21o] Anonymous. October 2021. Information Processing Letters, 171(??):??, October 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2021:S

[Ano21p]  Anonymous.  September

2021. Information Processing Letters, 170(??):??, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. Information Processing Letters, 175(??):??, April 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022.

Information Processing Letters, 177(??):??, August 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2022:EBa

[Ano22c]  Anonymous.  Editorial

Board. Information Processing Letters, 173(??):??, January 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001162.

Anonymous:2022:EBb

[Ano22d]  Anonymous.  Editorial

Board. Information Processing Letters, 174(??):??, March 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001320.

Anonymous:2022:EBc

[Ano22e]  Anonymous.  Editorial

Board. Information Processing Letters, 175(??):??, April 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001526.

Anonymous:2022:EBd

[Ano22f]  Anonymous.  Editorial

Board. Information Processing Letters, 176(??):??, June 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000151.

Anonymous:2022:EBe

[Ano22g]  Anonymous.  Editorial

Board. Information Processing Letters, 177(??):??, August 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000369.

Anonymous:2022:EBf

[Ano22h]  Anonymous.  Editorial

Board. Information Processing Letters, 178(??):??, November 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000667.

Anonymous:2022:Ja

[Ano22i] Anonymous. January 2022.

Information Processing Letters, 173(??):??, January 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2022:Jb

[Ano22j] Anonymous. June 2022. Information Processing Letters, 176(??):??, June 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2022:M

[Ano22k] Anonymous. March 2022. Information Processing Letters, 174(??):??, March 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2022:N

[Ano22l] Anonymous. November 2022.

Information Processing Letters, 178(??):??, November 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2023:A

[Ano23a] Anonymous. August 2023.

Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2023:EBa

[Ano23b]  Anonymous.  Editorial

Board. Information Processing Letters, 179(??):??, January 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000813.

Anonymous:2023:EBb

[Ano23c]  Anonymous.  Editorial

Board. Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022001053.

Anonymous:2023:EBc

[Ano23d]  Anonymous.  Editorial

Board. Information Processing Letters, 181(??):??, March 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000145.

Anonymous:2023:EBd

[Ano23e]  Anonymous.  Editorial

Board. Information Processing Letters, 182(??):??, August 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000558.

Anonymous:2023:F

[Ano23f] Anonymous. February 2023.

Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2023:J

[Ano23g] Anonymous. January 2023.

Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2023:Ma

[Ano23h] Anonymous. March 2023. Information Processing Letters, 181(??):??, March 2023. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2024:A

[Ano24a] Anonymous. August 2024.

Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2024:EBa

[Ano24b]  Anonymous.  Editorial

Board. Information Processing Letters, 183(??):??, January 2024. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000844.

Anonymous:2024:EBb

[Ano24c]  Anonymous.  Editorial

Board. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023001035.

Anonymous:2024:EBc

[Ano24d]  Anonymous.  Editorial

Board. Information Processing Letters, 185(??):??, March 2024. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000073.

Anonymous:2024:EBd

[Ano24e]  Anonymous.  Editorial

Board. Information Processing Letters, 186(??):??, August 2024. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000371.

Anonymous:2024:F

[Ano24f] Anonymous. February 2024. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2024:Ja

[Ano24g] Anonymous. January 2024.

Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2024:M

[Ano24h] Anonymous. March 2024. Information Processing Letters, 185(??):??, March 2024. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2025:EBa

[Ano25a]  Anonymous.  Editorial

Board. Information Processing Letters, 187(??):??, January 2025. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000577.

Anonymous:2025:EBb

[Ano25b]  Anonymous.  Editorial

Board. Information Processing Letters, 188(??):??, February 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000760.

Anonymous:2025:F

[Ano25c] Anonymous. February 2025. Information Processing Letters, 188(??):??, February 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2025:J

[Ano25d] Anonymous. January 2025. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alemany-Puig:2022:MPL

[APEiC22] Lluís Alemany-Puig, Juan Luis Esteban, and Ramon Ferrer i Cancho. Minimum projective linearizations of trees in linear time. Information Processing Letters, 174(??):??, March 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001198.

Alemany-Puig:2024:MLA

[APEiC24] Lluís Alemany-Puig, Juan Luis Esteban, and Ramon Ferrer i Cancho. The maximum linear arrangement problem for trees under projectivity and planarity. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000431.

Antony:2025:ASC

[APS25] Dhanyamol Antony, Sagartanu Pal, and R. B. Sandeep. Algorithms for subgraph complementation to some classes of graphs. Information Processing Letters, 188(??):??, February 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000607.

Atanasiu:2022:EEC

[APZT22] Adrian Atanasiu, Ghajendran Poovanandran, Abdalhadi Abu Zeyneh, and Wen Chean Teh. Erasure and error correcting ability of Parikh matrices. Information Processing Letters, 175(??):??, April 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001381.

Andrianarivelo:2022:MTP

[AR22] Nirina Andrianarivelo and Pierre Réty. Modular termination of prefix-constrained term rewrite systems. Information Processing Letters, 174(??):??, March 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001228.

Alvarez-Rebollar:2022:RPS

[ÁRCLM^{+}22] J. L. Álvarez-Rebollar, J. Cravioto-Lagos, N. Marín, E. Solís-Villarreal, and J. Urrutia. Representing point sets on the plane as permutations. Information Processing Letters, 175(??):??, April 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001435.

Abam:2021:GSP

[AS21] Mohammad Ali Abam and Mohammad Javad Rezaei Seraji. Geodesic spanners for points in R3 amid axis-parallel boxes. Information Processing Letters, 166(??):??, February 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301502.

Acosta:2024:SEC

[AT24] Nidia Obscura Acosta and Alexandru I. Tomescu. Simplicity in Eulerian circuits: Uniqueness and safety. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000649.

Ailon:2021:CCA

[AY21] Nir Ailon and Gal Yehuda.

The complexity of computing (almost) orthogonal matrices with ϵ-copies of the Fourier transform. Information Processing Letters, 165(??):??, January 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301113.

Bae:2022:FCE

[Bae22] Sang Won Bae. Faster counting empty convex polygons in a planar point set. Information Processing Letters, 175(??):??, April 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001368.

Beyersdorff:2021:SPQ

[BB21] Olaf Beyersdorff and Joshua Blinkhorn. A simple proof of QBF hardness. Information Processing Letters, 168(??):??, June 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000077.

Banik:2022:GSU

[BBBMS22] Aritra Banik, Bhaswar B.

Bhattacharya, Sujoy Bhore, and Leonardo Martínez-Sandoval. Geometric systems of unbiased representatives. Information Processing Letters, 176(??):??, June 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001472.

Becher:2021:EBS

[BC21] Verónica Becher and Lucas Cortés. Extending de Bruijn sequences to larger alphabets. Information Processing Letters, 168(??):??, June 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301721.

Bazgan:2020:GPT

[BCD20] Cristina Bazgan, Janka Chlebíková, and Clément Dallard. Graphs without a partition into two proportionally dense subgraphs. Information Processing Letters, 155(??):??, March 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301607.

Blazevic:2024:ATM

[BCEM24] Mislav Blazević, Stefan Canzar, Khaled Elbassioni, and Domagoj Matijević. Anti Tai mapping for unordered labeled trees. Information Processing Letters, 185 (??):??, March 2024. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000972.

Blazej:2023:PKT

[BCK^{+}23a] Václav Blažej, Pratibha Choudhary, Dušan Knop, Jan Matyáš Křišťan, Ondřej Suchý, and Tomáš Valla. Polynomial kernels for tracking shortest paths. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000722.

Bosch-Calvo:2023:IKF

[BCK23b] Miguel Bosch-Calvo and Steven Kelk. An improved kernel for the flip distance problem on simple convex polygons. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000248.

Berczi:2023:ARR

[BCKP23] Kristóf Bérczi, Karthekeyan Chandrasekaran, Tamás Király, and Aditya Pillai. Analyzing Residual Random Greedy for monotone submodular maximization. Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000977.

Baril:2021:CSP

[BCKV21] Jean-Luc Baril, Giulio Cerbai, Carine Khalil, and Vincent Vajnovszki. Catalan and Schröder permutations sortable by two restricted stacks. Information Processing Letters, 171(??):??, October 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000533.

Brunelli:2021:CPO

[BCV21] Filippo Brunelli, Pierluigi Crescenzi, and Laurent Viennot. On computing Pareto optimal paths in weighted time-dependent networks. Information Processing Letters, 168(??):??, June 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020301733.

Bereg:2025:CSL

[BDBK^{+}25] Sergey Bereg, José-Miguel Díaz-Báñez, Alina Kasiuk, Miguel-Angel Pérez-Cutiño, and Fabio Rodríguez. Covering segments on a line with drones. Information Processing Letters, 188(??): ??, February 2025. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902400070X.

Baril:2021:HTC

[BDH21] Ambroise Baril, Riccardo Dondi, and Mohammad Mehdi Hosseinzadeh. Hardness and tractability of the γ-complete subgraph problem. Information Processing Letters, 169(??):??, August 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000193.

Brosse:2024:HIW

[BDK^{+}24] Caroline Brosse, Oscar Defrain, Kazuhiro Kurita, Vincent Limouzy, Takeaki Uno, and Kunihiro Wasa. On the hardness of inclusionwise minimal separators enumeration. Information Processing Letters, 185(??):??, March 2024. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023001126.

Bednarczyk:2021:SEC

[Bed21] Bartosz Bednarczyk. Statistical EL is ExpTimecomplete. Information Processing Letters, 169(??):??, August 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000272.

Boyar:2020:RDO

[BEL20] Joan Boyar, Faith Ellen, and Kim S. Larsen. Randomized distributed online algorithms against adaptive offline adversaries. Information Processing Letters, 161(??):??, September 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300600.

Bhattacharya:2022:MMC

[BFJ22] Anup Bhattacharya, Yoav Freund, and Ragesh Jaiswal. On the k-means/median cost function. Information Processing Letters, 177(??):??, August 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000096.

Bertrand:2023:SWU

[BFM23] Nathalie Bertrand, Hugo Francon, and Nicolas Markey. Synchronizing words under LTL constraints. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000352.

Berard:2022:CRR

[BH22a] Béatrice Bérard and Serge Haddad. Corrigendum to “Revisiting reachability in polynomial interrupt timed automata” [Information Processing Letters 174 (2022) 106208]. Information Processing Letters, 175(??):??, April 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001460.

See [BH22b].

Berard:2022:RRP

[BH22b] Béatrice Bérard and Serge Haddad. Revisiting reachability in Polynomial Interrupt Timed Automata. Information Processing Letters, 174(??):??, March 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902100123X.

See corrigendum [BH22a].

Bhateja:2022:DAT

[Bha22]  Puneet Bhateja.  Deter-

mining asynchronous test equivalence for probabilistic processes. Information Processing Letters, 177(??): ??, August 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000266.

Brettell:2022:LCF

[BHMP22] Nick Brettell, Jake Horsfield, Andrea Munaro, and Daniël Paulusma. List kcolouring P_{t}-free graphs: a mim-width perspective. Information Processing Letters, 173(??):??, January 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000831.

Biedl:2020:SRS

[Bie20]  Therese Biedl.  Segment

representations with small resolution. Information Processing Letters, 153(??): ??, January 2020. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019301516.

Biedl:2022:HSN

[Bie22]  Therese Biedl.  Horton–

Strahler number, rooted pathwidth and upward drawings of trees. Information Processing Letters, 175 (??):??, April 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001459.

Bannai:2023:LBP

[BIK23] Hideo Bannai, Tomohiro I., and Dominik Köppl. Longest bordered and periodic subsequences. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000418.

Bannai:2021:LPO

[BIM21] Hideo Bannai, Shunsuke Inenaga, and Neerja Mhaskar. Longest previous overlapping factor array. Information Processing Letters, 168(??):??, June 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000119.

Bansal:2023:NTL

[BKK23] Nikhil Bansal, John Kuszmaul, and William Kuszmaul. A nearly tight lower bound for the d-dimensional cow-path problem. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000327.

Baruch:2023:GBW

[BKS23] Gilad Baruch, Shmuel T. Klein, and Dana Shapira. Guided blocks WOM codes. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019022000692.

Blikstad:2020:LCS

[Bli20] Joakim Blikstad. On the longest common subsequence of Thue–Morse words. Information Processing Letters, 164(??):??, December 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020301071.

Biniaz:2020:BMH

[BMS20] Ahmad Biniaz, Anil Maheshwari, and Michiel Smid. Bottleneck matchings and Hamiltonian cycles in higherorder Gabriel graphs. Information Processing Letters, 153(??):??, January 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019301528.

Barth:2022:PAD

[BMWW22] Dominique Barth, Thierry Mautor, Dimitri Watel, and Marc-Antoine Weisser. A polynomial algorithm for deciding the validity of an electrical distribution tree. Information Processing Letters, 176(??):??, June 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000060.

Belgi:2022:PAA

[BN22] Amir Belgi and Zeev Nutov. A polylogarithmic approximation algorithm for 2edge-connected dominating set. Information Processing Letters, 173(??):??, January 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000909.

Bodwin:2022:NDP

[Bod22]  Greg Bodwin.  A note

on distance-preserving graph sparsification. Information Processing Letters, 174 (??):??, March 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001204.

Brand:2022:NAT

[Bra22] Cornelius Brand. A note on algebraic techniques for subgraph detection. Information Processing Letters, 176(??):??, June 2022.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021001575.

Bistarelli:2021:KEC

[BRS21] Stefano Bistarelli, Fabio Rossi, and Francesco Santini. Kruskal with embedded Csemirings to solve MST problems with partially-ordered costs. Information Processing Letters, 169(??):??, August 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000211.

Billstein:2021:NIG

[BS21] Andreas Billstein and Rainer Schrader. A note on integral generalized flows in directed partial 2-trees. Information Processing Letters, 172(??):??, December 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000624.

Biswas:2023:MLS

[BS23] Aniruddha Biswas and Palash Sarkar. On the “majority is least stable” conjecture. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000527.

Buss:2024:RRE

[BY24] Sam Buss and Emre Yolcu.

Regular resolution effectively simulates resolution. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902400019X.

Couto:2022:SGT

[CCJS22] Fernanda Couto, Luís Felipe I. Cunha, Daniel Juventude, and Leandro Santiago. Strategies for generating tree spanners: Algorithms, heuristics and optimal graph classes. Information Processing Letters, 177(??):??, August 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000229.

Cicerone:2021:EGP

[CDDN21] Serafino Cicerone, Mattia D’Emidio, Gabriele Di Stefano, and Alfredo Navarra. On the effectiveness of the genetic paradigm for polygonization. Information Processing Letters, 171(??): ??, October 2021. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000491.

Ciccarelli:2023:NSP

[CDP23] Felice Ciccarelli, Miriam Di Ianni, and Giancarlo Palumbo. A note on the satisfactory partition problem: Constant size requirement. Information Processing Letters, 179(??):??, January 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000497.

Colomboni:2025:IUC

[CEP25] Roberto Colomboni, Emmanuel Esposito, and Andrea Paudice. An improved uniform convergence bound with fat-shattering dimension. Information Processing Letters, 188(??):??, February 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000693.

Chakraborti:2021:IRU

[CFHH21]  Debsoumya  Chakraborti,

Alan Frieze, Simi Haber, and Mihir Hasabnis. Isomorphism for random kuniform hypergraphs. Information Processing Letters, 166(??):??, February 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301265.

Choudhary:2023:IKT

[CGG^{+}23a] Pratibha Choudhary, Michael T.

Goodrich, Siddharth Gupta, Hadi Khodabandeh, Pedro Matias, and Venkatesh Raman. Improved kernels for tracking paths. Information Processing Letters, 181(??):??, March 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000030.

Cimatti:2023:GER

[CGG^{+}23b] Alessandro Cimatti, Luca Geatti, Nicola Gigante, Angelo Montanari, and Stefano Tonetta. GR(1) is equivalent to R(1). Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902200076X.

Chavrimootoo:2025:DGG

[Cha25] Michael C. Chavrimootoo.

Defying gravity and gadget numerosity: the complexity of the Hanano Puzzle and beyond. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000504.

Chen:2023:WPG

[Che23] Wei Chen. Weakest preconditioned goto axiom. Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000862.

Chillara:2020:SIL

[Chi20] Suryajith Chillara. Slightly improved lower bounds for homogeneous formulas of bounded depth and bounded individual degree. Information Processing Letters, 156(??):??, April 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019301838.

Chen:2021:TEL

[CHTW21] Xujin Chen, Xiaodong Hu, Zhongzheng Tang, and Chenhao Wang. Tight efficiency lower bounds for strategy-proof mechanisms in two-opposite-facility location game. Information Processing Letters, 168 (??):??, June 2021. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000120.

Correia:2020:CMF

[CIM20] Alexandre Correia, Juliano Iyoda, and Alexandre Mota. Combining model finder and genetic programming into a general purpose automatic program synthesizer. Information Processing Letters, 154(??):??, February 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301498.

Chakraborty:2023:ST

[CK23] Debrup Chakraborty and Samir Kundu. On the security of TrCBC. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000771.

Czumaj:2023:PTP

[CL23] Artur Czumaj and Andrzej Lingas. On parallel time in population protocols. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000710.

Chen:2025:AAI

[CLWX25] Jie Chen, Yi-Ping Liang, Cai-Xia Wang, and Shou-Jun Xu. Algorithmic aspects of P_{k}-isolation in graphs and extremal graphs for a P_{3}-isolation bound. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000516.

Csikos:2022:OAM

[CM22] Mónika Csikós and Nabil H. Mustafa. Optimal approximations made easy. Information Processing Letters, 176(??):??, June 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000072.

Cazaux:2020:HOG

[CR20] Bastien Cazaux and Eric Rivals. Hierarchical overlap graph. Information Processing Letters, 155(??): ??, March 2020. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301450.

Cummings:2023:MA

[CSS23] Robert Cummings, Jeffrey Shallit, and Paul Staadecker. Mesosome avoidance. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000485.

Chen:2022:CCO

[CST22] Da-Ren Chen, Min-Zheng Shieh, and Shi-Chun Tsai. The complexity of comparing optimal solutions. Information Processing Letters, 177(??):??, August 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000230.

Cifuentes:2023:CSS

[CST23] Santiago Cifuentes, Francisco J. Soulignac, and Pablo Terlisky. Complexity of solving a system of difference constraints with variables restricted to a finite set. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000212.

Castiglioni:2020:RLE

[CT20] Valentina Castiglioni and Simone Tini. Raiders of the lost equivalence: Probabilistic branching bisimilarity. Information Processing Letters, 159–160(??):??, July 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902030034X.

Cui:2020:DFR

[CWW20] Tingting Cui, Wei Wang, and Meiqin Wang. Distinguisher on full-round compression function of GOST R. Information Processing Letters, 156(??):??, April 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301851.

Chen:2021:CCG

[CX21] Jie Chen and Shou-Jun Xu. A characterization of 3-γ-critical graphs which are not bicritical. Information Processing Letters, 166(??):??, February 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301496.

Djukanovic:2021:SAC

[DBRB21] Marko Djukanovic, Christoph Berger, Günther R. Raidl, and Christian Blum. An A^{⋆} search algorithm for the constrained longest common subsequence problem. Information Processing Letters, 166(??):??, February 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301289.

Diskin:2023:HLP

[DE23] Sahar Diskin and Dor Elboim. Heavy and light paths and Hamilton cycles. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902300039X.

Deng:2022:CD

[Den22] Shichuan Deng. On clustering with discounts. Information Processing Letters, 177(??):??, August 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000291.

Datta:2020:LTD

[DFL^{+}20] A. K. Datta, P. Ferragina, L. Larmore, L. Pagli, and G. Prencipe. Linear time distributed swap edge algorithms. Information Processing Letters, 161(??):??, September 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300661.

Dimos:2023:EBP

[DFLS23] Sotirios Dimos, Dimitris Fotakis, Thanasis Lianeas, and Kyriakos Sergis. Escaping Braess’s paradox through approximate Caratheodory’s theorem. Information Processing Letters, 179(??):??, January 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000461.

Duque:2021:CNC

[DFMHVHT21] Frank Duque, Ruy Fabila-Monroy, César Hernández-Vélez, and Carlos Hidalgo-Toscano. Counting the number of crossings in geometric graphs. Information Processing Letters, 165(??):??, January 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301150.

Du:2022:CTS

[DFW22] Yusong Du, Baoying Fan, and Baodian Wei. A constant-time sampling algorithm for binary Gaussian distribution over the integers. Information Processing Letters, 176(??):??, June 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000035.

Dillencourt:2023:SCB

[DG23] Michael Dillencourt and Michael T. Goodrich. Simplified Chernoff bounds with powers-of-two probabilities. Information Processing Letters, 182(??):??, August 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000406.

Devanny:2021:CAS

[DGI21] William E. Devanny, Michael T.

Goodrich, and Sandy Irani. A competitive analysis for the Start-Gap algorithm for online memory wear leveling. Information Processing Letters, 166(??):??, February 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301290.

Dillencourt:2025:LPC

[DGM25] Michael Dillencourt, Michael T.

Goodrich, and Michael Mitzenmacher. Leveraging parameterized Chernoff bounds for simplified algorithm analyses. Information Processing Letters, 187(??):??, January 2025. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000462.

Dondi:2023:CDS

[DH23] Riccardo Dondi and Danny Hermelin. Computing the k densest subgraphs of a graph. Information Processing Letters, 179(??):??, January 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000734.

Dinh:2024:TIN

[DH24] Jérémi Do Dinh and Alexandros Hollender. Tight inapproximability of Nash equilibria in public goods games. Information Processing Letters, 186(??):??, August 2024. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000164.

Davis:2022:PCZ

[DHP^{+}22] Ben Davis, Hamed Hatami, William Pires, Ran Tao, and Hamza Usmani. On publiccoin zero-error randomized communication complexity. Information Processing Letters, 178(??):??, November 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000503.

Dahlberg:2022:CVM

[DHW22] Axel Dahlberg, Jonas Helsen, and Stephanie Wehner. The complexity of the vertexminor problem. Information Processing Letters, 175 (??):??, April 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902100137X.

DeCarufel:2021:TPT

[DK21] Jean-Lou De Carufel and Tanvir Kaykobad. Transforming plane triangulations by simultaneous diagonal flips. Information Processing Letters, 170 (??):??, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902100034X.

Dahiya:2024:LTF

[DKMS24] Yogesh Dahiya, Vignesh K., Meena Mahajan, and Karteek Sreenivasaiah. Linear threshold functions in decision lists, decision trees, and depth-2 circuits. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000613.

Doka:2020:DPR

[DKP^{+}20] K. Doka, A. Kosmatopoulos, A. Papadopoulos, S. Sioutas, K. Tsichlas, and D. Tsoumakos. Dynamic planar range skyline queries in log logarithmic expected time. Information Processing Letters, 162(??):??, October 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300776.

Dundua:2024:PUF

[DKS24] Besik Dundua, Ioane Kapanadze, and Helmut Seidl. Prenex universal first-order safety properties. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000188.

Das:2023:RLP

[DLN^{+}23] Tapas Das, Tuomo Lehtilä, Soumen Nandi, Sagnik Sen, and D. K. Supraja. On radio k-labeling of the power of the infinite path. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000297.

Dobraunig:2020:PFO

[DMM20] Christoph Dobraunig, Florian Mendel, and Bart Mennink. Practical forgeries for ORANGE. Information Processing Letters, 159–160(??):??, July 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902030048X.

Dybizbanski:2020:SCD

[DNS20] Janusz Dybizbański, Anna Nenca, and Andrzej Szepietowski. Signed coloring of 2-dimensional grids. Information Processing Letters, 156(??):??, April 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300053.

Doerr:2021:RAE

[Doe21]  Benjamin Doerr.  Run-

time analysis of evolutionary algorithms via symmetry arguments. Information Processing Letters, 166(??):??, February 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301514.

Doyen:2025:TCA

[Doy25]  Laurent Doyen.  Top-

down complementation of automata on finite trees. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000292.

Dragan:2020:EAS

[Dra20] Feodor F. Dragan. An eccentricity 2-approximating spanning tree of a chordal graph is computable in linear time. Information Processing Letters, 154(??):??, February 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301565.

Dybizbanski:2021:HCP

[DS21] Janusz Dybizbański and Andrzej Szepietowski. Hamiltonian cycles and paths in hypercubes with disjoint faulty edges. Information Processing Letters, 172(??):??, December 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000727.

DellErba:2024:SFG

[DSTZ24] Daniele Dell’Erba, Sven Schewe, Qiyi Tang, and Tansholpan Zhanabekova. Semantic flowers for goodfor-games and deterministic automata. Information Processing Letters, 185 (??):??, March 2024. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023001114.

Durr:2023:IBR

[Dür23]  Anita Dürr.  Improved

bounds for rectangular monotone Min-Plus Product and applications. Information Processing Letters, 181(??): ??, March 2023. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000017.

Eom:2024:LTA

[EA24] Taekang Eom and Hee-Kap Ahn. A linear-time algorithm for the center problem in weighted cycle graphs. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000255.

ElAtik:2021:CPM

[EAE21] Abd El Fattah A. El Atik, A. W. Aboutahoun, and A. Elsaid. Correct proof of the main result in “The number of spanning trees of a class of self-similar fractal models” by Ma and Yao. Information Processing Letters, 170(??):??, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000314.

Englert:2021:LBC

[EHL^{+}21] Matthias Englert, Piotr Hofman, Sławomir Lasota, Ranko Lazić, Jérôme Leroux, and Juliusz Straszyński. A lower bound for the coverability problem in acyclic pushdown VAS. Information Processing Letters, 167 (??):??, April 2021. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301666.

Echenim:2022:EUS

[EIP22] Mnacho Echenim, Radu Iosif, and Nicolas Peltier. Entailment is undecidable for symbolic heap separation logic formulæ with non-established inductive rules. Information Processing Letters, 173(??):??, January 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000843.

Eriguchi:2020:SSL

[EK20] Reo Eriguchi and Noboru Kunihiro. Strong security of linear ramp secret sharing schemes with general access structures. Information Processing Letters, 164(??):??, December 2020.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301058.

Engelfriet:2021:CMS

[Eng21]  Joost Engelfriet.  Com-

putability by monadic secondorder logic. Information Processing Letters, 167(??): ??, April 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301617.

Elaroussi:2023:PEA

[ENRV23] Mohammed Elaroussi, Lhouari Nourine, Mohammed Said Radjef, and Simon Vilmin. On the preferred extensions of argumentation frameworks: Bijections with naive sets. Information Processing Letters, 181(??):??, March 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022001119.

Echenim:2023:URS

[EP23] Mnacho Echenim and Nicolas Peltier. An undecidability result for Separation Logic with theory reasoning. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000029.

Feghali:2023:NMC

[Feg23]  Carl Feghali.  A note

on matching-cut in P_{t}-free graphs. Information Processing Letters, 179(??):??, January 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000515.

Fomin:2024:FAS

[FGIK24] Fedor V. Fomin, Petr A. Golovach, Tanmay Inamdar, and Tomohiro Koana. FPT approximation and subexponential algorithms for covering few or many edges. Information Processing Letters, 185(??):??, March 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000012.

Fiorenza:2023:PRV

[FGS23] Dario Fiorenza, Daniele Gorla, and Ivano Salvo. Polynomial recognition of vulnerable multi-commodities. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000394.

Foucaud:2021:CAI

[FHL21] Florent Foucaud, Hervé Hocquard, and Dimitri Lajou. Complexity and algorithms for injective edgecoloring in graphs. Information Processing Letters, 170(??):??, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000351.

Fu:2023:QCS

[FHL^{+}23] Jianling Fu, Cheng-Chao Huang, Yong Li, Jingyi Mei, Ming Xu, and Lijun Zhang. Quantitative controller synthesis for consumption Markov decision processes. Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000990.

Fiorino:2022:NCT

[Fio22] Guido Fiorino. A non-clausal tableau calculus for Min-Sat. Information Processing Letters, 173(??):??, January 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902100082X.

Farhana:2025:SUH

[FK25] Tsuri Farhana and Matthew J. Katz. Spanners under the Hausdorff and Fréchet distances. Information Processing Letters, 187(??):??, January 2025. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000437.

Fujimori:2020:FAM

[FKMS20] Yusei Fujimori, Yasushi Kawase, Tomomi Matsui, and Akiyoshi Shioura. A fast algorithm for multiprocessor speed-scaling problem minimizing completion time and energy consumption. Information Processing Letters, 162(??):??, October 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300788.

Fryganiotis:2023:NNC

[FPP23] Nikolaos Fryganiotis, Symeon Papavassiliou, and Christos Pelekis. A note on the network coloring game: a randomized distributed (Δ + 1)- coloring algorithm. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000285.

Franzle:2020:EDR

[FQSW20] Martin Fränzle, Karin Quaas, Mahsa Shirmohammadi, and James Worrell. Effective definability of the reachability relation in timed automata. Information Processing Letters, 153(??):??, January 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301541.

Fritsch:2021:OGE

[Fri21] Robin Fritsch. Online graph exploration on trees, unicyclic graphs and cactus graphs. Information Processing Letters, 168(??):??, June 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000107.

Fukunaga:2022:TLH

[FRRT22] Takuro Fukunaga, R. Ravi, Oleksandr Rudenko, and Ziye Tang. Two-level hub Steiner trees. Information Processing Letters, 174 (??):??, March 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001241.

Fleischer:2020:NBA

[FRS20] Lukas Fleischer, Samin Riasat, and Jeffrey Shallit. New bounds on antipowers in words. Information Processing Letters, 164 (??):??, December 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020301083.

Fernstrom:2020:CAA

[FS20] Finn Fernstrøm and Teresa Anna Steiner. A constant approximation algorithm for the uniform a priori capacitated vehicle routing problem with unit demands. Information Processing Letters, 159–160(??):??, July 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300478.

Fujito:2023:NAD

[Fuj23] Toshihiro Fujito. A note on approximations of directed edge dominating set. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000606.

Gabric:2024:RUB

[Gab24]  Daniel Gabric.  Ranking

and unranking bordered and unbordered words. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000959.

Gillani:2021:QNB

[GB21] Iqra Altaf Gillani and Amitabha Bagchi. A queueing network-based distributed Laplacian solver for directed graphs. Information Processing Letters, 166(??):??, February 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301277.

Gonzalez:2020:CGC

[GGSdS20] Lucía M. González, Luciano N. Grippo, Martín D. Safe, and Vinicius F. dos Santos. Covering graphs with convex sets and partitioning graphs into convex sets. Information Processing Letters, 158(??):??, June 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300314.

Ghodrati:2025:LUB

[GH25] Amir Hossein Ghodrati and Mohammad Ali Hosseinzadeh. Lower and upper bounds on graph communicabilities. Information Processing Letters, 187(??): ??, January 2025. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000310.

Gildea:2020:MQR

[GHKY20] Joe Gildea, Holly Hamilton, Abidin Kaya, and Bahattin Yildiz. Modified quadratic residue constructions and new extremal binary self-dual codes of lengths 64, 66 and 68. Information Processing Letters, 157(??):??, May 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300144.

Gima:2025:ISL

[GHN^{+}25] Tatsuya Gima, Tesshu Hanaka, Kohei Noro, Hirotaka Ono, and Yota Otachi. An improved spectral lower bound of treewidth. Information Processing Letters, 188(??): ??, February 2025. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000668.

Giannella:2021:IRE

[Gia21] Chris R. Giannella. Instability results for Euclidean distance, nearest neighbor search on high dimensional Gaussian data. Information Processing Letters, 169(??):??, August 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000296.

Genc:2020:SPO

[GIR20] Ziya Alper Genç, Vincenzo Iovino, and Alfredo Rial. “The simplest protocol for oblivious transfer” revisited. Information Processing Letters, 161 (??):??, September 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300624.

Goyal:2023:TFA

[GJ23] Dishant Goyal and Ragesh Jaiswal. Tight FPT approximation for Socially Fair Clustering. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000261.

Gottlieb:2022:NUP

[GK22] Lee-Ad Gottlieb and Aryeh Kontorovich. Non-uniform packings. Information Processing Letters, 174(??):??, March 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000946.

Grabowski:2023:SEH

[GK23] Szymon Grabowski and Dominik Köppl. Space-efficient Huffman codes revisited. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001902200031X.

Georgiou:2023:OCS

[GKL^{+}23] Konstantinos Georgiou, Evangelos Kranakis, Nikos Leonardos, Aris Pagourtzis, and Ioannis Papaioannou. Optimal circle search despite the presence of faulty robots. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000340.

Guingona:2023:CAP

[GKNS23] Vincent Guingona, Alexei Kolesnikov, Julianne Nierwinski, and Avery Schweitzer. Comparing approximate and probabilistic differential privacy parameters. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000236.

Gomez:2022:IEB

[GKP22] Ana I. Gomez, Markus Kiderlen, and Florian Pausinger. Improved entropy bounds for parity filtered self-timed ring based random number generators. Information Processing Letters, 174 (??):??, March 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001277.

Gluch:2025:RTL

[GL25] Grzegorz Głuch and Krzysztof Loryś. 4/3 rectangle tiling lower bound. Information Processing Letters, 188(??): ??, February 2025. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902400053X.

Guo:2023:AAV

[GLW23] Lifeng Guo, Changhong Lu, and Guanlin Wu. Approximation algorithms for a virtual machine allocation problem with finite types. Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000965.

Gaikwad:2022:GMD

[GM22] Ajinkya Gaikwad and Soumen Maity. Globally minimal defensive alliances. Information Processing Letters, 177(??):??, August 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000102.

Guelev:2024:ECS

[GM24] Dimitar P. Guelev and Ben Moszkowski. Expressive completeness by separation for discrete time interval temporal logic with expanding modalities. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000103.

Goeminne:2020:CEW

[Goe20]  Aline Goeminne.  Con-

strained existence of weak subgame perfect equilibria in multiplayer Büchi games. Information Processing Letters, 163(??):??, November 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300831.

Grandoni:2022:RAE

[GOR^{+}22] Fabrizio Grandoni, Rafail Ostrovsky, Yuval Rabani, Leonard J. Schulman, and Rakesh Venkat. A refined approximation for Euclidean k-means. Information Processing Letters, 176 (??):??, June 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000084.

Gil-Pons:2024:FHF

[GPWM24] Reynaldo Gil-Pons, Max Ward, and Loïc Miller. Finding (s, d)-hypernetworks in F -hypergraphs is NPhard. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000765.

Galesi:2024:VCN

[GRZ24] Nicola Galesi, Fariba Ranjbar, and Michele Zito. Vertex-connectivity for node failure identification in Boolean Network Tomography. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000935.

Gabric:2021:BPP

[GS21] Daniel Gabric and Jeffrey Shallit. Borders, palindrome prefixes, and square prefixes. Information Processing Letters, 165(??):??, January 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020301149.

Gadducci:2022:DRL

[GS22] Fabio Gadducci and Francesco Santini. Distributivity and residuation for lexicographic orders. Information Processing Letters, 177(??):??, August 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001902200028X.

Giakkoupis:2025:LMA

[GTZ25] George Giakkoupis, Volker Turau, and Isabella Ziccardi. Luby’s MIS algorithms made self-stabilizing. Information Processing Letters, 188(??):??, February 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000619.

Gao:2021:TBN

[GW21] Wei Gao and Weifan Wang. Tight binding number bound for P_{≥3}-factor uniform graphs. Information Processing Letters, 172(??):??, December 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000776.

Hong:2024:GFP

[HBL24] Haibo Hong, Shi Bai, and Fenghao Liu. The group factorization problem in finite groups of Lie type. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000140.

Hayamizu:2020:RRC

[HHMM20] Momoko Hayamizu, Katharina T. Huber, Vincent Moulton, and Yukihiro Murakami. Recognizing and realizing cactus metrics. Information Processing Letters, 157(??):??, May 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902030003X.

Hadzilacos:2022:RCR

[HHT22] Vassos Hadzilacos, Xing Hu, and Sam Toueg. Randomized consensus with regular registers. Information Processing Letters, 174 (??):??, March 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000880.

Hentschel:2023:EPS

[HHT23] Brian Hentschel, Peter J. Haas, and Yuanyuan Tian. Exact PPS sampling with bounded sample size. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001902300025X.

Huang:2022:OEI

[HJHZ22] Yan Huang, Yan Jin, Zhi Hu, and Fangguo Zhang. Optimizing the evaluation of lisogenous curve for isogenybased cryptography. Information Processing Letters, 178(??):??, November 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000588.

Hershkowitz:2020:RGB

[HK20] D. Ellis Hershkowitz and Gregory Kehne. Reverse greedy is bad for k-center. Information Processing Letters, 158(??):??, June 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300284.

Hansen:2021:TBR

[HKP21] Jakob Cetti Hansen, Adam Husted Kjelstrøm, and Andreas Pavlogiannis. Tight bounds for reachability problems on one-counter and pushdown systems. Information Processing Letters, 171 (??):??, October 2021. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000508.

Hershkowitz:2021:ORH

[HKR21] D. Ellis Hershkowitz, Gregory Kehne, and R. Ravi. An optimal rounding for half-integral weighted minimum strongly connected spanning subgraph. Information Processing Letters, 167(??):??, April 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902030154X.

Hinrichsen:2020:LPF

[HLS20] Erica G. Hinrichsen, Valeria A. Leoni, and Martín D. Safe. Labelled packing functions in graphs. Information Processing Letters, 154(??):??, February 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019301462.

Hasan:2024:RPG

[HMR24] Md. Manzurul Hasan, Debajyoti Mondal, and Md. Saidur Rahman. Relating planar graph drawings to planar satisfiability problems. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000893.

Henning:2020:AHR

[HPP20] Michael A. Henning, Saikat Pal, and D. Pradhan. Algorithm and hardness results on hop domination in graphs. Information Processing Letters, 153(??): ??, January 2020. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019301553.

Hakim:2022:NRP

[HPR22] Sheikh Azizul Hakim, Bishal Basak Papan, and Md. Saidur Rahman. New results on pairwise compatibility graphs. Information Processing Letters, 178(??):??, November 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000412.

Hochbaum:2020:AAC

[HR20] Dorit S. Hochbaum and Xu Rao. Approximation algorithms for connected maximum coverage problem for the discovery of mutated driver pathways in cancer. Information Processing Letters, 158(??):??, June 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300272.

Harada:2020:RSL

[HS20] Masaaki Harada and Ken Saito. Remark on subcodes of linear complementary dual codes. Information Processing Letters, 159–160(??):??, July 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300508.

Hoppenot:2024:RAI

[HS24a] Pierre Hoppenot and Zoltán Szigeti. On reversing arcs to improve arc-connectivity. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000777.

Horsch:2024:SCP

[HS24b] Florian Hörsch and Zoltán Szigeti. Steiner connectivity problems in hypergraphs. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000716.

Hartman:2025:RSU

[HS25] Samuel Hartman and Jonathan P. Sorenson. Reducing the space used by the sieve of Eratosthenes when factoring. Information Processing Letters, 188(??):??, February 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902400067X.

Halldorsson:2021:CIV

[HT21] Magnús M. Halldórsson and Tigran Tonoyan. Computing inductive vertex orderings. Information Processing Letters, 172(??):??, December 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000740.

Huang:2023:PD

[Hua23] Ming-Deh A. Huang. On product decomposition. Information Processing Letters, 181(??):??, March 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022001016.

Huber:2025:NPF

[HvIMS25] Katharina T. Huber, Leo van Iersel, Vincent Moulton, and Guillaume E. Scholz. Is this network proper forest-based? Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000309.

Halman:2021:RAR

[HW21] Nir Halman and Shmuel Wimer. Resource allocation in rooted trees subject to sum constraints and nonlinear cost functions. Information Processing Letters, 170(??):??, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000284.

Huang:2022:DLB

[HW22] Xiaomin Huang and Chenhao Wang. Discrete load balancing on complete bipartite graphs. Information Processing Letters, 175 (??):??, April 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021001393.

Huang:2020:GCC

[HYZ^{+}20] Meijuan Huang, Bo Yang, Yi Zhao, Xin Wang, Yanwei Zhou, and Zhe Xia. A generic construction of CCA-secure deterministic encryption. Information Processing Letters, 154(??):??, February 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019301486.

Imana:2023:DOM

[ID23] José L. Imaña and Siemen Dhooghe. Domain-oriented masked bit-parallel finitefield multiplier against sidechannel attacks. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000388.

Indzhev:2022:CUA

[IK22] Emil Indzhev and Stefan Kiefer. On complementing unambiguous automata and graphs with many cliques and cocliques. Information Processing Letters, 177 (??):??, August 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000278.

Ikenmeyer:2022:NVC

[IS22] Christian Ikenmeyer and Abhiroop Sanyal. A note on VNP-completeness and border complexity. Information Processing Letters, 176(??):??, June 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001587.

Ishizuka:2021:CFC

[Ish21] Takashi Ishizuka. On the complexity of finding a Caristi’s fixed point. Information Processing Letters, 170(??):??, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000338.

Ismailov:2024:AES

[Ism24] Vugar E. Ismailov. Approximation error of single hidden layer neural networks with fixed weights. Information Processing Letters, 185(??):??, March 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023001102.

Ito:2025:ORA

[Ito25] Fuki Ito. Optimal randomized algorithms of weaklybalanced multi-branching AND–OR trees. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000425.

Jamshidpour:2020:SAD

[JA20] Sadegh Jamshidpour and Zahra Ahmadian. Security analysis of a dynamic threshold secret sharing scheme using linear subspace method. Information Processing Letters, 163 (??):??, November 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300818.

Jacques:2021:CNS

[Jac21] Fabien Jacques. On the chromatic numbers of signed triangular and hexagonal grids. Information Processing Letters, 172(??):??, December 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000715.

Jaiswal:2020:NRB

[Jai20] Ragesh Jaiswal. A note on the relation between XOR and Selective XOR lemmas. Information Processing Letters, 163(??):??, November 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300983.

Jena:2022:DTD

[JJD22] Sangram K. Jena, Ramesh K.

Jallu, and Gautam K. Das. On d-distance m-tuple (l, r)- domination in graphs. Information Processing Letters, 174(??):??, March 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000934.

Jobson:2021:MCG

[JKL21] Adam S. Jobson, André E.

Kézdy, and Jenő Lehel. Minimal 2-connected graphs satisfying the even cut condition. Information Processing Letters, 167(??):??, April 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301678.

Jowhari:2025:ARU

[JN25] Hossein Jowhari and Shamisa Nematollahi. Airports and railways with unsplittable demand. Information Processing Letters, 188(??):??, February 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000681.

Jain:2022:SIC

[JPV22] Vishesh Jain, Huy Tuan Pham, and Thuy-Duong Vuong. Spectral independence, coupling, and the spectral gap of the Glauber dynamics. Information Processing Letters, 177(??):??, August 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000254.

Jansen:2020:NIG

[JR20] Klaus Jansen and Lars Rohwedder. A note on the integrality gap of the configuration LP for restricted Santa Claus. Information Processing Letters, 164 (??):??, December 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301125.

Jowhari:2023:MPC

[JR23] Hossein Jowhari and Mohsen Rezapour. Monochromatic partitioning of colored points by lines. Information Processing Letters, 182(??):??, August 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000455.

Jukna:2020:SCE

[JS20] Stasys Jukna and Hannes Seiwert. Sorting can exponentially speed up pure dynamic programming. Information Processing Letters, 159–160(??):??, July 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300491.

Jacob:2024:LDD

[JWZ24] Ashwin Jacob, Michał Włodarczyk, and Meirav Zehavi. Long directed detours: Reduction to 2-disjoint paths. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000218.

Januszewski:2022:PBI

[JZ22] Janusz Januszewski and Lukasz Zielonka. Packing batches of items into a single bin. Information Processing Letters, 174 (??):??, March 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021001113.

Januszewski:2023:PBC

[JZ23] Janusz Januszewski and Lukasz Zielonka. Packing batches of cubes into a single bin. Information Processing Letters, 180(??): ??, February 2023. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000941.

Kamiyama:2021:EFM

[Kam21] Naoyuki Kamiyama. The envy-free matching problem with pairwise preferences. Information Processing Letters, 172(??):??, December 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000739.

Kamiyama:2023:OPA

[Kam23] Naoyuki Kamiyama. On optimization problems in acyclic hypergraphs. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000339.

Kim:2021:SFI

[KC21] Sung-Hwan Kim and Hwan-Gue Cho. Simpler FMindex for parameterized string matching. Information Processing Letters, 165(??):??, January 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301137.

Kiyomi:2021:LCS

[KHO21] Masashi Kiyomi, Takashi Horiyama, and Yota Otachi. Longest common subsequence in sublinear space. Information Processing Letters, 168(??):??, June 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902030171X.

Kisek:2021:CAS

[KK21] Anja Kisek and Sandi Klavzar. Correcting the algorithm for the secure domination number of cographs by Jha, Pradhan, and Banerjee. Information Processing Letters, 172(??):??, December 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000703.

Kim:2023:OPP

[KKNS23] Youngho Kim, Munseong Kang, Joong Chae Na, and Jeong Seop Sim. Orderpreserving pattern matching with scaling. Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000904.

Kowaluk:2020:SAN

[KL20] Mirosław Kowaluk and Andrzej Lingas. A simple approach to nondecreasing paths. Information Processing Letters, 162(??):??, October 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902030079X.

Kelk:2023:CGI

[KLM23] Steven Kelk, Simone Linz, and Ruben Meuwese. Cyclic generators and an improved linear kernel for the rooted subtree prune and regraft distance. Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902200093X.

Klug:2024:CMS

[Klu24] Nikolas Klug. Computing minimal solutions to the ring loading problem. Information Processing Letters, 185(??):??, March 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023001096.

Konstantinova:2021:SCG

[KM21] Elena V. Konstantinova and Alexey N. Medvedev. Small cycles, generalized prisms and Hamiltonian cycles in the bubble-sort graph. Information Processing Letters, 168(??):??, June 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000089.

Kashyop:2020:LED

[KN20] Manas Jyoti Kashyop and N. S. Narayanaswamy. Lazy or eager dynamic matching may not be fast. Information Processing Letters, 162(??):??, October 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300697.

Knop:2021:LLS

[Kno21] Dusan Knop. Local linear set on graphs with bounded twin cover number. Information Processing Letters, 170(??):??, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000326.

Kostolanyi:2023:SDQ

[Kos23] Peter Kostolányi. On some decidability questions concerning supports of rational series. Information Processing Letters, 179(??): ??, January 2023. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000473.

Kostolanyi:2025:FPS

[Kos25]  Peter Kostolányi.  Free

products of semigroups and monoids with a deterministic context-free word problem. Information Processing Letters, 188(??):??, February 2025. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000711.

Kaznatcheev:2021:WAC

[KP21] Artem Kaznatcheev and Prakash Panangaden. Weighted automata are compact and actively learnable. Information Processing Letters, 171(??):??, October 2021.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000478.

Kitaev:2024:STO

[KP24] Sergey Kitaev and Artem Pyatkin. On semi-transitive orientability of split graphs. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000789.

Komarath:2020:CDH

[KS20] Balagopal Komarath and Nitin Saurabh. On the complexity of detecting hazards. Information Processing Letters, 162(??):??, October 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300673.

Krauthgamer:2022:SSN

[KS22] Robert Krauthgamer and Shay Sapir. Smoothness of Schatten norms and slidingwindow matrix streams. Information Processing Letters, 177(??):??, August 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000114.

Keikha:2023:VVM

[KS23] Vahideh Keikha and Maria Saumell. On Voronoi visibility maps of 1.5D terrains with multiple viewpoints. Information Processing Letters, 181(??):??, March 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000054.

Laber:2024:CCS

[Lab24] Eduardo Sany Laber. The computational complexity of some explainable clustering problems. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000807.

Lingg:2024:LPN

[LdOOW24] Jonas Lingg, Mateus de Oliveira Oliveira, and Petra Wolf. Learning from positive and negative examples: New proof for binary alphabets. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000704.

Levin:2022:RAP

[Lev22]  Asaf Levin.  Robust al-

gorithms for preemptive scheduling on uniform machines of non-increasing job sizes. Information Processing Letters, 174(??):??, March 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001265.

Luckow:2020:CCL

[LF20] Max-Jonathan Luckow and Till Fluschnik. On the computational complexity of length- and neighborhoodconstrained path problems. Information Processing Letters, 156(??):??, April 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301966.

Li:2023:BSU

[LG23] Shuguang Li and Zhichao Geng. Bicriteria scheduling on an unbounded parallelbatch machine for minimizing makespan and maximum cost. Information Processing Letters, 180(??): ??, February 2023. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022001004.

Liu:2023:NIR

[Liu23]  Quanquan C. Liu.  A

note on improved results for one round distributed clique listing. Information Processing Letters, 181 (??):??, March 2023. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022001120.

Luy:2022:CES

[LKC22] Erkam Luy, Zekeriya Y. Karatas, and Olcay Ciftci. Comment on “An enhanced and secured RSA public cryptosystem algorithm using Chinese remainder theorem (ESRPKC)”. Information Processing Letters, 177(??):??, August 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000205.

Li:2024:FLP

[LL24a] Xiaowei Li and Xiwen Lu. The facility location problem with maximum distance constraint. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001902300090X.

Lienardy:2024:WOU

[LL24b] Jean Liénardy and Frédéric Lafitte. A weakness in OCB3 used with short nonces allowing for a break of authenticity and confidentiality. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000479.

Lan:2025:HCB

[LL25] Ting Lan and Huazhong Lü.

Hamiltonian cycles of balanced hypercube with disjoint faulty edges. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000486.

Li:2021:SSC

[LLC21] Jing Li, Xujing Li, and Eddie Cheng. Super spanning connectivity of splitstar networks. Information Processing Letters, 166(??): ??, February 2021. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301241.

Li:2023:SEC

[LLLW23] Xiangwen Li, Yangfan Li, Jian-Bo Lv, and Tao Wang. Strong edge-colorings of sparse graphs with 3Δ − 1 colors. Information Processing Letters, 179(??):??, January 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000709.

Lafourcade:2020:ABI

[LLP20] Pascal Lafourcade and Marius Lombard-Platet. About blockchain interoperability. Information Processing Letters, 161(??):??, September 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300636.

Lyon:2022:IDP

[LM22] Merritt Lyon and Hosam Mahmoud. Insertion depth in power-weight trees. Information Processing Letters, 176(??):??, June 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021001423.

Lozin:2020:IDV

[LMMZ20] Vadim Lozin, Dmitriy Malyshev, Raffaele Mosca, and Viktor Zamaraev. Independent domination versus weighted independent domination. Information Processing Letters, 156(??):??, April 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300016.

Luca:2022:GSL

[LMO^{+}22] V. T. F. Luca, N. Marín, F. S. Oliveira, A. Ramírez-Vigueras, O. Solé-Pi, J. L. Szwarcfiter, and J. Urrutia. Grid straight-line embeddings of trees with a minimum number of bends per path. Information Processing Letters, 174(??):??, March 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001253.

Lopez:2025:RLT

[LMT25] Théodore Lopez, Benjamin Monmege, and Jean-Marc Talbot. Regular D-length: a tool for improved prefixstable forward Ramsey factorisations. Information Processing Letters, 187(??): ??, January 2025. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000279.

Louza:2020:SAC

[Lou20] Felipe A. Louza. A simple algorithm for computing the document array. Information Processing Letters, 154(??):??, February 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001901930170X.

Loukides:2022:APS

[LP22] Grigorios Loukides and Solon P. Pissis. All-pairs suffix/prefix in optimal time using Aho–Corasick space. Information Processing Letters, 178(??):??, November 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000321.

Li:2020:AOP

[LPT20] Yanjun Li, Jie Peng, and Chik How Tan. An answer to an open problem of mesnager on bent functions. Information Processing Letters, 161(??):??, September 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300612.

Lucero:2023:NCL

[LS23] Jorge C. Lucero and Sławek Staworko. A note on the class of languages generated by F -systems over regular languages. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000400.

Lin:2025:NRP

[LT25] Jing-You Lin and Shi-Chun Tsai. A note on the krestriction problem. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000450.

Liu:2023:CRO

[LTT23] Zi-Yuan Liu, Yi-Fan Tseng, and Raylin Tso. Cryptanalysis of a round optimal lattice-based multisignature scheme. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000078.

Lee:2023:SHA

[LW23] Euiwoong Lee and Pengxiang Wang. Strong hardness of approximation for tree transversals. Information Processing Letters, 181(??):??, March 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022001090.

Li:2023:SLB

[LXZW23] Kang Li, Fengjun Xiao, Bingpeng Zhou, and Jinming Wen. A sharper lower bound on Rankin’s constant. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000224.

Lu:2022:OSM

[LZG22] Xiwen Lu, Kejun Zhao, and Manzhan Gu. Online single-machine scheduling to minimize the linear combination of makespans of two agents. Information Processing Letters, 173(??):??, January 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000788.

Mackie:2024:RFM

[Mac24] Maximilien Mackie. Recursionfree modular arithmetic in the lambda-calculus. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000510.

Manurangsi:2021:LDH

[Man21]  Pasin Manurangsi.  Lin-

ear discrepancy is Π_{2}-hard to approximate. Information Processing Letters, 172(??):??, December 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902100079X.

Manurangsi:2024:NHC

[Man24] Pasin Manurangsi. A note on hardness of computing recursive teaching dimension. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000728.

Manurangsi:2025:ILB

[Man25] Pasin Manurangsi. Improved lower bound for differentially private facility location. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000322.

Mastakas:2021:DRT

[Mas21] Konstantinos Mastakas. Drawing a rooted tree as a rooted y-monotone minimum spanning tree. Information Processing Letters, 166(??):??, February 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020301228.

Morrison:2020:OSB

[MG20] Benjamin Morrison and Adam Groce. Oracle separations between quantum and non-interactive zeroknowledge classes. Information Processing Letters, 154(??):??, February 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019019301474.

Mili:2021:DD

[Mil21]  Ali Mili.  Differentiators

and detectors. Information Processing Letters, 169 (??):??, August 2021. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000259.

Mirzanezhad:2024:ANN

[Mir24] Majid Mirzanezhad. On approximate near-neighbors search under the (continuous) Fréchet distance in higher dimensions. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000480.

Muller:2020:NCM

[MK20] Julian Müller and Sven Kosub. A note on the complexity of manipulating weighted Schulze voting. Information Processing Letters, 162(??):??, October 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300764.

Mayank:2020:PTS

[MM20] Jaishree Mayank and Arijit Mondal. Polynomial time schedulability test for periodic non-preemptive 2task system. Information Processing Letters, 154(??): ??, February 2020. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301504.

Ma:2020:NSM

[MMCH20] Tianlong Ma, Yaping Mao, Eddie Cheng, and Ping Han. A note on the strong matching preclusion problem for data center networks. Information Processing Letters, 164(??):??, December 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300946.

Meng:2020:TCS

[MMHX20] Keju Meng, Fuyou Miao, Wenchao Huang, and Yan Xiong. Threshold changeable secret sharing with secure secret reconstruction. Information Processing Letters, 157(??):??, May 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300156.

Mor:2020:LSS

[MMS20] Baruch Mor, Gur Mosheiov, and Dana Shapira. Lot scheduling on a single machine to minimize the (weighted) number of tardy orders. Information Processing Letters, 164(??):??, December 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902030096X.

Molter:2022:CFT

[Mol22] Hendrik Molter. The complexity of finding temporal separators under waiting time constraints. Information Processing Letters, 175(??):??, April 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001447.

Mooij:2022:STC

[Moo22]  Arjan J. Mooij.  Static

type checking without downcast operator. Information Processing Letters, 178 (??):??, November 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000424.

Mafort:2020:VDS

[MP20] Rodrigo Lamblet Mafort and Fábio Protti. Vector domination in splitindifference graphs. Information Processing Letters, 155(??):??, March 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301826.

Minaud:2023:GCH

[MP23] Brice Minaud and Charalampos Papamanthou. Generalized cuckoo hashing with a stash, revisited. Information Processing Letters, 181(??):??, March 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022001132.

Martin:2022:HPQ

[MPS22] Barnaby Martin, Daniël Paulusma, and Siani Smith. Hard problems that quickly become very easy. Information Processing Letters, 174(??):??, March 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001289.

Mishra:2025:DPC

[MRD25] Pawan K. Mishra, S. V. Rao, and Gautam K. Das. Dispersion problem on a convex polygon. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000280.

Mucha:2020:IAF

[MS20a] Marcin Mucha and Marcin Smulewicz. Improved approximation for fractionally subadditive network design. Information Processing Letters, 154(??):??, February 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301449.

Mulzer:2020:CPC

[MS20b] Wolfgang Mulzer and Natalia Shenkman. A constructive proof of a concentration bound for realvalued random variables. Information Processing Letters, 158(??):??, June 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300296.

Maneth:2023:DOE

[MS23] Sebastian Maneth and Helmut Seidl. Deciding origin equivalence of weakly selfnesting macro tree transducers. Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000898.

Maji:2024:DMT

[MS24a] Sukanya Maji and Sanjib Sadhu. Discrete and mixed two-center problems for line segments. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000947.

Maneth:2024:CPT

[MS24b] Sebastian Maneth and Helmut Seidl. Checking in polynomial time whether or not a regular tree language is deterministic top-down. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000923.

Manurangsi:2023:MBM

[MSHS23] Pasin Manurangsi, Erel Segal-Halevi, and Warut Suksompong. On maximum bipartite matching with separation. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000315.

Mitsunobu:2024:WCA

[MSS24] Takuto Mitsunobu, Reiji Suda, and Vorapong Suppakitpaisarn. Worst-case analysis of LPT scheduling on a small number of nonidentical processors. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000674.

Morita:2024:FPD

[MSYY24] Kohei Morita, Shinya Shiroshita, Yutaro Yamaguchi, and Yu Yokoi. Fast primaldual update against local weight update in linear assignment problem and its application. Information Processing Letters, 183(??):

??, January 2024. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000753.

Mishiba:2020:QEC

[MT20] Shohei Mishiba and Yasuhiko Takenaga. QUIXO is EXPTIME-complete. Information Processing Letters, 162(??):??, October 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902030082X.

Ma:2021:SCP

[MT21] Yanger Ma and Tony Tan.

A simple combinatorial proof for the small model property of two-variable logic. Information Processing Letters, 170(??):??, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000363.

Mikawa:2023:ELT

[MT23] Kenji Mikawa and Ken Tanaka. Efficient lineartime ranking and unranking of derangements. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902200045X.

Misselbeck-Wessel:2023:MEM

[MW23] Daniel Misselbeck-Wessel. Maximal elements with minimal logic. Information Processing Letters, 182(??): ??, August 2023. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000467.

Mieno:2022:PTS

[MWN^{+}22] Takuya Mieno, Kiichi Watanabe, Yuto Nakashima, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Palindromic trees for a sliding window and its applications. Information Processing Letters, 173(??):??, January 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000892.

Ma:2018:NST

[MY18] Fei Ma and Bing Yao. The number of spanning trees of a class of self-similar fractal models. Information Processing Letters, 136 (??):64–69, August 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901830084X.

See corrected proof [EAE21].

Nies:2024:WAG

[NS24] André Nies and Frank Stephan. Word automatic groups of nilpotency class

2. Information Processing Letters, 183(??):??, January 2024. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000698.

Nutov:2025:AAM

[Nut25]  Zeev Nutov.  A 22k-

approximation algorithm for minimum power k edge disjoint st-paths. Information Processing Letters, 188(??):??, February 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000620.

Ordanel:2021:PTA

[OFA21] Ivy Ordanel, Proceso Fernandez, and Henry Adorna. A polynomial time algorithm for the 2-Poset Cover Problem. Information Processing Letters, 169(??):??, August 2021. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902100020X.

Ohsaka:2021:FPP

[Ohs21]  Naoto Ohsaka.  A fully

polynomial parameterized algorithm for counting the number of reachable vertices in a digraph. Information Processing Letters, 171(??):??, October 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000521.

Omar:2023:CMT

[OPD23] Satyam Omar, Sahadeo Padhye, and Dhananjoy Dey. Cryptanalysis of multivariate threshold ring signature schemes. Information Processing Letters, 181 (??):??, March 2023. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022001144.

Oropeza:2021:RCT

[OT21] Marcos Oropeza and Csaba D. Tóth. Reconstruction of the crossing type of a point set from the compatible exchange graph of noncrossing spanning trees. Information Processing Letters, 170(??):??, September 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000302.

Park:2023:DRM

[PB23] Moonju Park and Hyeongboo Baek. Determining rate monotonic schedulability of real-time periodic tasks using continued fractions. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000539.

Pai:2020:WEC

[PCC20] Kung-Jui Pai, Ruay-Shiung Chang, and Jou-Ming Chang. A well-equalized 3-CIST partition of alternating group graphs. Information Processing Letters, 155(??):??, March 2020. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301577.

Paskin-Cherniavsky:2020:CAU

[PCO20] Anat Paskin-Cherniavsky and Ruxandra F. Olimid. On cryptographic anonymity and unpredictability in secret sharing. Information Processing Letters, 161 (??):??, September 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300521.

Poureidi:2023:ARR

[PF23] Abolfazl Poureidi and Jafar Fathali. Algorithmic results in Roman dominating functions on graphs. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000066.

Pham:2023:NAS

[PH23] Canh V. Pham and Dung T. K. Ha. A note for approximating the submodular cover problem over integer lattice with low adaptive and query complexities. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000364.

Patawar:2023:DDS

[PK23] Maithilee Patawar and Kalpesh Kapoor. Density of distinct squares in non-primitive words. Information Processing Letters, 182(??):??, August 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000108.

Park:2024:SWC

[PK24] Je Hong Park and Woo-Hwan Kim. Security weakness of a certificate-based proxy signature scheme for IIoT environments. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000492.

Poureidi:2022:CNB

[Pou22] Abolfazl Poureidi. On computing the number of (BC- )subtrees, eccentric subtree number, and global and local means of trees. Information Processing Letters, 178(??):??, November 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902200059X.

Pattanayak:2024:DTH

[PP24] Debasish Pattanayak and Andrzej Pelc. Deterministic treasure hunt and rendezvous in arbitrary connected graphs. Information Processing Letters, 185 (??):??, March 2024. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000984.

Panagiotou:2024:EIA

[PR24] Konstantinos Panagiotou and Simon Reisser. The effect of iterativity on adversarial opinion forming. Information Processing Letters, 185(??):??, March 2024.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000960.

Panda:2024:CCS

[PRM24] B. S. Panda, Soumyashree Rana, and Sounaka Mishra. On the complexity of cosecure dominating set problem. Information Processing Letters, 185(??):??, March 2024. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023001060.

Potechin:2020:LWA

[PS20] Aaron Potechin and Jeffrey Shallit. Lengths of words accepted by nondeterministic finite automata. Information Processing Letters, 162(??):??, October 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300806.

Pudlak:2022:MPU

[Pud22] Pavel Pudlák. On matrices potentially useful for tree codes. Information Processing Letters, 174(??): ??, March 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000958.

Paluch:2021:SCA

[PW21] Katarzyna Paluch and Mateusz Wasylkiewicz. A simple combinatorial algorithm for restricted 2matchings in subcubic graphs

— via half-edges. Information Processing Letters, 171(??):??, October 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000612.

Polak:2024:LAM

[PZ24] Adam Polak and Maksym Zub. Learning-augmented maximum flow. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000176.

Qiu:2024:ICI

[QW24] Guoliang Qiu and Jiaheng Wang. Inapproximability of counting independent sets in linear hypergraphs. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000911.

Rabinovich:2022:GBR

[Rab22] Yuri Rabinovich. A generalization of the Blind Rotating Table game. Information Processing Letters, 176(??):??, June 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021001484.

Ray:2024:TNA

[Ray24] Arka Ray. There is no AP-TAS for 2-dimensional vector bin packing: Revisited. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001902300073X.

Rahmati:2020:RDS

[RE20] Zahed Rahmati and Fatemeh Emami. RAC drawings in subcubic area. Information Processing Letters, 159–160(??):??, July 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300326.

Rauch:2023:ERG

[RR23] Johannes Rauch and Dieter Rautenbach. Efficiently recognizing graphs with equal independence and annihilation numbers. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000303.

Raman:2020:CKE

[RRS20] Venkatesh Raman, M. S. Ramanujan, and Saket Saurabh. A characterization of König– Egerváry graphs with extendable vertex covers. Information Processing Letters, 161(??):??, September 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902030051X.

Rauch:2025:CFC

[RRS25] Johannes Rauch, Dieter Rautenbach, and Uéverton S. Souza. On conflict-free cuts: Algorithms and complexity. Information Processing Letters, 187(??):??, January 2025. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000334.

Raman:2023:PCM

[RSRM23] Remi Raman, Shahin John J.

S., Subashini R., and Subhasree Methirumangalath. On the parameterized complexity of the Maximum Exposure Problem. Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000953.

Ravi:2020:ARL

[RT20a] Peruvemba Sundaram Ravi and Levent Tunçel. Approximation ratio of LD algorithm for multi-processor scheduling and the Coffman– Sethi conjecture. Information Processing Letters, 159–160(??):??, July 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300466.

Raynal:2020:MEF

[RT20b] Michel Raynal and Gadi Taubenfeld. Mutual exclusion in fully anonymous shared memory systems. Information Processing Letters, 158(??):??, June 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300259.

See corrigendum [RT23].

Raynal:2023:CME

[RT23] Michel Raynal and Gadi Taubenfeld. Corrigendum to “Mutual exclusion in fully anonymous shared memory systems” [Inf. Process. Lett. 158 (2020) 105938]. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000618.

Rukavicka:2020:UBN

[Ruk20]  Josef Rukavicka.  Upper

bound for the number of closed and privileged words. Information Processing Letters, 156(??):??, April 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300041.

Rescigno:2023:BAG

[RV23] Adele A. Rescigno and Ugo Vaccaro. Bounds and algorithms for generalized superimposed codes. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902300008X.

Sahbi:2022:NSL

[Sah22a] Rafik Sahbi. New sharp lower bound for the quorum coloring number of trees. Information Processing Letters, 178(??):??, November 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000540.

Sahin:2022:NNE

[Sah22b] Bünyamin Sahin. New network entropy: the domination entropy of graphs. Information Processing Letters, 174(??):??, March 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021001101.

Sakharov:2020:ARE

[Sak20] Alexander Sakharov. Annotated regular expressions and input-driven languages. Information Processing Letters, 159–160(??):??, July 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300454.

Saxena:2021:ZTA

[Sax21] Sanjeev Saxena. Zone theorem for arrangements in dimension three. Information Processing Letters, 172(??):??, December 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000764.

Smyth:2022:SDE

[SC22] Ben Smyth and Michael R. Clarkson. Surveying definitions of election verifiability. Information Processing Letters, 177(??):??, August 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000242.

Schilling:2021:RCR

[Sch21] Judith Schilling. Results and conjectures on the role of the uniform distribution in the coupon collector’s problem with group drawings. Information Processing Letters, 169(??):??, August 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000260.

Scheffler:2024:RLT

[Sch24] Robert Scheffler. Recognizing LBFS trees of bipartite graphs. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000139.

Sellke:2024:SIS

[Sel24]  Mark Sellke.  On size-

independent sample complexity of ReLU networks. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0020019024000127.

Severin:2020:ACN

[Sev20] Daniel Severín. On the additive chromatic number of several families of graphs. Information Processing Letters, 158(??):??, June 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300247.

Shang:2020:LDN

[Sha20] Yilun Shang. Longest distance of a non-uniform dispersion process on the infinite line. Information Processing Letters, 164(??): ??, December 2020. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300958.

Shallit:2021:RAM

[Sha21] Jeffrey Shallit. Robbins and Ardila meet Berstel. Information Processing Letters, 167(??):??, April 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001902030168X.

Shang:2023:LPH

[Sha23] Yilun Shang. Long paths in heterogeneous random subgraphs of graphs with large minimum degree. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000443.

Shibutani:2022:IPI

[SI22a] Kyoji Shibutani and Tetsu Iwata. On the (im)possibility of improving the round diffusion of generalized Feistel structures. Information Processing Letters, 174 (??):??, March 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001125.

Shinagawa:2022:QAS

[SI22b] Kazuo Shinagawa and Tetsu Iwata. Quantum attacks on sum of even-Mansour pseudorandom functions. Information Processing Letters, 173(??):??, January 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000879.

Singh:2023:IPS

[Sin23] Abhishek Singh. On the intractability of preemptive single-machine job scheduling with release times, deadlines, and family setup times. Information Processing Letters, 179(??):??, January 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001902200062X.

Sabir:2021:FTH

[SM21] Eminjan Sabir and Jixiang Meng. Fault-tolerant Hamiltonicity of hypercubes with faulty subcubes. Information Processing Letters, 172(??):??, December 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000752.

Smyth:2020:SGV

[Smy20]  Ben Smyth.  Surveying

global verifiability. Information Processing Letters, 163(??):??, November 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300879.

Sokol:2020:DPM

[Sok20] Dina Sokol. 2-dimensional palindromes with k mismatches. Information Processing Letters, 164(??):??, December 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902030106X.

Singh:2020:IBB

[SP20] Sonika Singh and Sahadeo Padhye. Identity based blind signature scheme over NTRU lattices. Information Processing Letters, 155 (??):??, March 2020. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301814.

Shubham:2022:ASA

[SPG22] Shubham, Surya Prakash, and Pramod Ganapathi. An algorithm for the sequence alignment with gap penalty problem using multiway divide-and-conquer and matrix transposition. Information Processing Letters, 173(??):??, January 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000818.

Schierreich:2022:WRB

[SS22] Simon Schierreich and Ondrej Suchý. Waypoint routing on bounded treewidth graphs. Information Processing Letters, 173(??):??, January 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000806.

Stankovic:2022:RMC

[Sta22] Aleksa Stanković. On regularity of Max-CSPs and Min-CSPs. Information Processing Letters, 176(??):??, June 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000011.

Stewart:2020:VNC

[Ste20]  Iain A. Stewart.  Vari-

ational networks of cubeconnected cycles are recursive cubes of rings. Information Processing Letters, 157(??):??, May 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300120.

Suksompong:2025:WFD

[Suk25] Warut Suksompong. Weighted fair division of indivisible items: a review. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000498.

Suppakitpaisarn:2022:TLB

[Sup22] Vorapong Suppakitpaisarn. Tight lower bound for average number of terms in optimal double-base number system using informationtheoretic tools. Information Processing Letters, 175 (??):??, April 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001411.

Sutra:2020:CEP

[Sut20] Pierre Sutra. On the correctness of Egalitarian Paxos. Information Processing Letters, 156(??):??, April 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001901930184X.

Tankus:2022:WWC

[Tan22] David Tankus. Weighted well-covered graphs without cycles of lengths 5, 6 and 7. Information Processing Letters, 174(??):??, March 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001046.

tenCate:2024:NEP

[tCFJL24] Balder ten Cate, Maurice Funk, Jean Christoph Jung, and Carsten Lutz. On the non-efficient PAC learnability of conjunctive queries. Information Processing Letters, 183(??):??, January 2024.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000741.

Tucker-Foltz:2023:TSC

[TF23] Jamie Tucker-Foltz. Thou shalt covet the average of thy neighbors’ cakes. Information Processing Letters, 180(??):??, February 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000989.

T:2025:CAR

[TGS25] Shabana A. T., Alphy George, and S. Sheerazuddin. Counter abstraction for regular open teams. Information Processing Letters, 188(??):??, February 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000632.

Tan:2024:NKR

[TP24] Chik How Tan and Theo Fanuela Prabowo. A new key recovery attack on a code-based signature from the Lyubashevsky framework. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000650.

Tsur:2020:FAC

[Tsu20a]  Dekel Tsur.  Faster al-

gorithms for cograph edge modification problems. Information Processing Letters, 158(??):??, June 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300338.

Tsur:2020:FAO

[Tsu20b] Dekel Tsur. An FPT algorithm for orthogonal buttons and scissors. Information Processing Letters, 163(??):??, November 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300843.

Tsur:2021:ADP

[Tsu21a]  Dekel Tsur.  Algorithms

for deletion problems on split graphs. Information Processing Letters, 167 (??):??, April 2021. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301538.

Tsur:2021:FPA

[Tsu21b]  Dekel Tsur.  Faster pa-

rameterized algorithm for Bicluster Editing. Information Processing Letters, 168(??):??, June 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000090.

Tsur:2021:KFE

[Tsu21c]  Dekel Tsur.  Kernel for

K_{t}-free edge deletion. Information Processing Letters, 167(??):??, April 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301691.

Tsur:2022:CDR

[Tsu22] Dekel Tsur. Cluster deletion revisited. Information Processing Letters, 173(??): ??, January 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000867.

Tsur:2023:FDAa

[Tsu23a] Dekel Tsur. Faster deterministic algorithm for Cactus Vertex Deletion. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000746.

Tsur:2023:FDAb

[Tsu23b] Dekel Tsur. Faster deterministic algorithm for Co-Path Set. Information Processing Letters, 180(??): ??, February 2023. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000928.

Tsur:2024:SKT

[Tsu24] Dekel Tsur. Smaller kernels for two vertex deletion problems. Information Processing Letters, 186 (??):??, August 2024. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000231.

Trobst:2023:RPB

[TV23] Thorben Tröbst and Vijay V.

Vazirani. A real polynomial for bipartite graph minimum weight perfect matchings. Information Processing Letters, 179(??):??, January 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000436.

Tian:2023:OCN

[TY23] Fangyu Tian and Yuxue Yin. The odd chromatic number of a toroidal graph is at most 9. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000273.

Tan:2025:RML

[TZZ25] Lixing Tan, Zhaohui Zhu, and Jinjin Zhang. Reduced meet over labellingbased semantics in abstract argumentation. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000449.

Utsumi:2024:EOB

[UNSI24] Shion Utsumi, Motoki Nakahashi, Kosei Sakamoto, and Takanori Isobe. Exploring the optimality of byte-wise permutations of a piccolotype block cipher. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000790.

Urschel:2021:TGP

[UW21] John C. Urschel and Jake Wellens. Testing gap kplanarity is NP-complete. Information Processing Letters, 169(??):??, August 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020301708.

Vahanwala:2024:SPC

[Vah24] Mihir Vahanwala. Skolem and positivity completeness of ergodic Markov chains. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000115.

vanBevern:2020:OSP

[vBS20] René van Bevern and Pavel V.

Smirnov. Optimal-size problem kernels for d-hitting set in linear time and space. Information Processing Letters, 163(??):??, November 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300855.

vanderHoog:2020:MAT

[vdHKL^{+}20] Ivor van der Hoog, Vahideh Keikha, Maarten Löffler, Ali Mohades, and Jérôme Urhausen. Maximum-area triangle in a convex polygon, revisited. Information Processing Letters, 161 (??):??, September 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300302.

Verma:2025:ICM

[VDPT25] Bhisham Dev Verma, Punit Pankaj Dubey, Rameshwar Pratap, and Manoj Thakur. Improving compressed matrix multiplication using control variate method. Information Processing Letters, 187(??):??, January 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000474.

Vigna:2020:POR

[Vig20] Sebastiano Vigna. On the probability of overlap of random subsequences of pseudorandom number generators. Information Processing Letters, 158(??):??, June 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300260.

vanIersel:2022:ARL

[vIKMN22] Leo van Iersel, Sjors Kole, Vincent Moulton, and Leonie Nipius. An algorithm for reconstructing level-2 phylogenetic networks from trinets. Information Processing Letters, 178(??):??, November 2022. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000576.

vanIersel:2023:PIC

[vIMM23] Leo van Iersel, Vincent Moulton, and Yukihiro Murakami. Polynomial invariants for cactuses. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000376.

Volkovich:2023:FNC

[Vol23] Ilya Volkovich. The final nail in the coffin of statisticallysecure obfuscator. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000091.

Verma:2020:GCS

[VP20] Shaily Verma and B. S.

Panda. Grundy coloring in some subclasses of bipartite graphs and their complements. Information Processing Letters, 163 (??):??, November 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300867.

Verma:2024:SCS

[VPD24] Bhisham Dev Verma, Rameshwar Pratap, and Punit Pankaj Dubey. Sparsifying count sketch. Information Processing Letters, 186(??):??, August 2024. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000206.

Verma:2024:UEI

[VPT24] Bhisham Dev Verma, Rameshwar Pratap, and Manoj Thakur. Unbiased estimation of inner product via higher order count sketch. Information Processing Letters, 183(??):??, January 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000509.

Wolf:2025:EDR

[WHN25] Rosa Wolf, Klaus Heeger, and André Nichterlein. Effective data reduction for strongly stable matching in very sparse graphs. Information Processing Letters, 188(??):??, February 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000644.

Wang:2021:RSC

[WL21] Longchun Wang and Qingguo Li. Representations of stably continuous semilattices by information systems and abstract bases. Information Processing Letters, 165(??):??, January 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902030123X.

Wei:2022:FRS

[WL22] Yao Wei and Zihui Liu. Further results on the second relative greedy weight of 3dimensional codes. Information Processing Letters, 178(??):??, November 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000552.

Wu:2023:PWB

[WL23] Hao Wu and Huan Long.

Probabilistic weak bisimulation and axiomatization for probabilistic models. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902300042X.

Wang:2020:OTC

[WQ20] Maoqun Wang and Jianguo Qian. An Ore-type condition for the existence of two disjoint cycles. Information Processing Letters, 159–160(??):??, July 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300442.

Wu:2024:RSM

[WTP24] Wei Wu, Liang Tang, and Andrea Pizzuti. Robust scheduling for minimizing maximum lateness on a serial-batch processing machine. Information Processing Letters, 186(??):??, August 2024. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000036.

Wu:2024:CMP

[WW24] Yu-Lun Wu and Hung-Lung Wang. Correcting matrix products over the ring of integers. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000267.

Wang:2024:AID

[WXC24] Shanshan Wang, Chenglong Xiao, and Emmanuel Casseau. Algorithms with improved delay for enumerating connected induced subgraphs of a large cardinality. Information Processing Letters, 183(??):??, January 2024. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000686.

Weimann:2020:IDP

[WY20] Oren Weimann and Raphael Yuster. Incremental distance products via faulty shortest paths. Information Processing Letters, 161 (??):??, September 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019020300648.

Wu:2024:BBS

[WYL24] Hao Wu, Qizhe Yang, and Huan Long. Branching bisimulation semantics for quantum processes. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002001902400022X.

Wei:2024:LER

[WYZ^{+}24] Qi Wei, Xiaolin Yao, Wenxin Zhang, Ruiyue Zhang, and Yonggong Ren. On-line exploration of rectangular cellular environments with a rectangular hole. Information Processing Letters, 185(??):??, March 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023001138.

Wu:2025:TBG

[WZ25] Haoxuan Wu and Jincheng Zhuang. Tighter bound for generalized multiple discrete logarithm problem via MDS matrix method. Information Processing Letters, 188(??):??, February 2025. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019024000656.

Wu:2022:NUB

[WZDZ22] Haoxuan Wu, Jincheng Zhuang, Qianheng Duan, and Yuqing Zhu. Nonuniform birthday problem revisited: Refined analysis and applications to discrete logarithms. Information Processing Letters, 175(??):??, April 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001902100140X.

Xiao:2020:ISD

[Xia20] Han Xiao. On ideal semicomplete digraphs. Information Processing Letters, 157(??):??, May 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301863.

Xiao:2022:SIP

[XK22] Mingyu Xiao and Shaowei Kou. A simple and improved parameterized algorithm for bicluster editing. Information Processing Letters, 174(??):??, March 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001083.

Xiao:2020:SRO

[XN20] Mingyu Xiao and Hiroshi Nagamochi. Some reduction operations to pairwise compatibility graphs. Information Processing Letters, 153(??):??, January 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019019301589.

Yang:2023:EEC

[Yan23]  Yuxing Yang.  Embed-

ded edge connectivity of kary n-cubes. Information Processing Letters, 180(??): ??, February 2023. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000850.

Yang:2024:SNN

[YHK24] Dar-Li Yang, Yung-Tsung Hou, and Wen-Hung Kuo. A short note on “A note on single-machine scheduling with job-dependent learning effects”. Information Processing Letters, 183(??): ??, January 2024. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000662.

Yan:2024:ACQ

[YKC24] Feifei Yan, Pinhui Ke, and Zuling Chang. The autocorrelation of a class of quaternary sequences of length pq with high complexity. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000243.

Yu:2022:APA

[YL22] Wei Yu and Zhaohui Liu. Approximation and polynomial algorithms for the data mule scheduling with handling time and time span constraints. Information Processing Letters, 178(??):??, November 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000564.

Yuan:2022:ULB

[YQL22] Jun Yuan, Huijuan Qiao, and Aixia Liu. The upper and lower bounds of R_{g}-conditional diagnosability of networks. Information Processing Letters, 176 (??):??, June 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000059.

Zamfirescu:2022:VDC

[Zam22] Carol T. Zamfirescu. Vertex degrees and 2-cuts in graphs with many Hamiltonian vertex-deleted subgraphs. Information Processing Letters, 174(??):??, March 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021001071.

Zhao:2023:CSC

[ZC23] Shu-Li Zhao and Jou-Ming Chang. Connectivity, super connectivity and generalized 3-connectivity of folded divide-and-swap cubes. Information Processing Letters, 182(??):??, August 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019023000200.

Zhao:2021:ICA

[ZCWW21] Zishen Zhao, Shiyao Chen, Meiqin Wang, and Wei Wang. Improved cubeattack-like cryptanalysis of reduced-round Ketje-Jr and Keccak-MAC. Information Processing Letters, 171(??): ??, October 2021. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000387.

Zeitz:2023:NHS

[Zei23] Tim Zeitz. NP-hardness of shortest path problems in networks with non-FIFO time-dependent travel times. Information Processing Letters, 179(??):??, January 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019022000448.

Zimmermann:2022:AML

[Zim22] Martin Zimmermann. Approximating the minimal lookahead needed to win infinite games. Information Processing Letters, 177 (??):??, August 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000217.

Zimmermann:2025:RPT

[Zim25] Martin Zimmermann. Robust probabilistic temporal logics. Information Processing Letters, 188(??): ??, February 2025. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000528.

Zheng:2024:NCG

[ZKP^{+}24] Lijing Zheng, Haibin Kan, Jie Peng, Yanjun Li, and Yanbin Zheng. A new class of generalized almost perfect nonlinear monomial functions. Information Processing Letters, 184(??):??, February 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019023000881.

Zhang:2024:SNC

[ZQKL24] Liyu Zhang, Mahmoud Quweider, Fitra Khan, and Hansheng Lei. Splitting NPcomplete sets infinitely. Information Processing Letters, 186(??):??, August 2024. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019024000024.

Zschoche:2022:FPA

[Zsc22] Philipp Zschoche. A faster parameterized algorithm for temporal matching. Information Processing Letters, 174(??):??, March 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002001902100096X.

Zou:2022:EVG

[ZWWC22] Meibiao Zou, Zhifeng Wang, Jianxin Wang, and Yixin Cao. End vertices of graph searches on bipartite graphs. Information Processing Letters, 173(??): ??, January 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000910.

Zhang:2020:OLS

[ZXH20] Yong Zhang, Jiayi Xian, and Menghu Huang. Online leasing strategy for depreciable equipment considering opportunity cost. Information Processing Letters, 162(??):??, October 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019020300685.

Zhu:2022:CIE

[ZXY^{+}22] Fei Zhu, Feihong Xu, Xu Yang, Xun Yi, and Alsharif Abuadbba.

Cryptanalysis and improvements of an efficient certificatebased proxy signature scheme for IIoT environments. Information Processing Letters, 173(??):??, January 2022. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019021000855.

Zhou:2023:PKE

[ZXZ^{+}23] Yanwei Zhou, Ran Xu, Wenzheng Zhang, Zhe Xia, Bo Yang, Chunxiang Gu, and Meijuan Huang. Publickey encryption scheme with optimal continuous leakage resilience. Information Processing Letters, 180(??): ??, February 2023. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022000758.

Zhao:2023:RLD

[ZY23] Zishen Zhao and Qing Ye.

Revisiting lower dimension lattice attacks on NTRU. Information Processing Letters, 181(??):??, March 2023. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019022001107.

Zhang:2021:RSO

[ZZ21] Meng Zhang and Yi Zhang.

Rank and select operations on a word. Information Processing Letters, 172(??):??, December 2021. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000636.

Zhang:2022:ATO

[ZZLC22] Yubai Zhang, Zhao Zhang, Zhaohui Liu, and Qirong Chen. An asymptotically tight online algorithm for m-Steiner Traveling Salesman Problem. Information Processing Letters, 174 (??):??, March 2022. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0020019021000922.